### v1

# A Two-Stage Fourth-Order Implicit Scheme for Stiff problems

Abstract
This paper presents a novel two-stage fourth-order implicit scheme designed to overcome the limitations of existing spatiotemporal coupling two-stage fourth-order methods, which have thus far been confined to explicit frameworks. In such frameworks, computational efficiency is severely constrained by the CFL condition when addressing stiff problems, and they fundamentally conflicts with spatiotemporal coupled methods whose core advantage lies in embedding stiff source terms. By fully leveraging temporal derivatives of physical variables within a rigorously derived compact mathematical framework, the scheme achieves fourth-order temporal accuracy in only two stages. Furthermore, a sufficient condition for A-stability is established through systematic theoretical and numerical investigation, and a Newton iterative procedure is provided to accelerate convergence. Extensive numerical experiments on classical stiff problems confirm the method’s effectiveness and competitiveness.
Key words: two-stage fourth-order; temporal derivatives; implicit scheme; A-stability; stiff problems; Newton iteration
1 Introduction
The generalized Riemann problem (GRP) fundamentally differs from the classical Riemann problem through its use of spatially varying initial data, which generates non-self-similar solutions with non-zero time derivatives. This critical advancement enables the natural incorporation of physical source terms and other essential effects (e.g., continuous entropy variation) into the temporal evolution, establishing GRP as an indispensable framework for capturing complex physical mechanisms. The simplest GRP form replaces piecewise constant with piecewise linear initial data in homogeneous systems, first explored by Kolgan [1] and further developed by van Leer [2, 3]. Toro and Titarev [4] later extended GRP to include source terms using the Lax-Wendroff procedure, forming the basis of arbitrarily high-order (ADER) schemes [5, 6] applicable to hydrodynamics, magnetohydrodynamics [7], reaction-diffusion [8], and stiff systems [9], though these linearized solvers struggle under extreme thermodynamic conditions. Ben-Artzi et al. [10, 11, 12] were the first to develop the GRP solver for compressible fluids, which can be considered as the nonlinear version of the ADER solver. By incorporating thermodynamic effects, the GRP solver can accurately resolve strong waves, making it valuable under extreme conditions [13]. The GRP solver has since been successfully applied to duct flow [14], reactive flow [15], shallow water equations [16, 17], multi-material and multiphase flow [18, 19, 20, 21], magnetohydrodynamics [22], relativistic hydrodynamics [23, 24, 25], blood flow modeling [26], and radiation hydrodynamics [27].
The inherent spatiotemporal coupling in fluid dynamics necessitates numerical algorithms that preserve such correlations, which conventional Riemann solvers fail to achieve as they only provide piecewise constant states for constructing instantaneous Cauchy fluxes. This limitation leads to error accumulation over time in flows with strong discontinuities. In contrast, generalized Riemann solvers (e.g., ADER or GRP) provide temporal derivatives of physical quantities, enabling construction of time-interval fluxes with inherent spatiotemporal coupling. These fluxes exhibit Lipschitz continuity with respect to boundary perturbations, allowing temporal effects to enhance spatial regularity. Consequently, errors from strong discontinuities can be directly controlled through the time step, enabling accurate long-term resolution while establishing consistency between mathematical weak solutions and physical integral balance laws [28, 29].
Runge-Kutta methods, originally developed by Carl Runge [30] in 1895 and Martin Wilhelm Kutta [31] in 1901, constitute a family of classical numerical techniques for solving initial-value problems of ordinary differential equations that comprise both explicit and implicit forms. The explicit variants, exemplified by the classical fourth-order RK4 scheme, achieve high-order accuracy through weighted averages of function values at multiple stages, offering implementation simplicity without requiring calculation of higher-order derivatives. However, their stability remains constrained by the CFL condition, rendering them inefficient for stiff problems. In contrast, Implicit Runge-Kutta methods proposed by Kuntzmann [32] and Butcher [33], such as Radau and Gauss-type schemes, compute slopes at future time steps by solving systems of nonlinear equations. Although computationally more expensive, they exhibit superior numerical stability, making them particularly suitable for stiff differential equations.
Recently, Li et al. developed a two-stage fourth-order time advancement method [34], which makes full use of the temporal derivatives of physical quantities, enabling the scheme to achieve fourth-order time accuracy in only two stages. This method successfully combines the simplicity advantages of Runge-Kutta-type methods with the spatiotemporally coupling properties of the GRP solver. Then they details the characteristics of this class of methods [35]: (i) Computational efficiency: Under the same conditions, its computational efficiency is twice that of the Runge-Kutta method, which requires four stages to achieve fourth-order time accuracy [36]; (ii) Compactness: Since the number of time advancement steps is reduced by half, the stencil size is also halved; the spatial-temporal coupled HWENO-type reconstruction [37, 38] can make the computational scheme more compact; (iii) Stability: It has been proven that its stability region is larger than that of Runge-Kutta methods [39]; (iv) Compatibility: The method can be easily combined with other methods, such as the GKS solver [40], the multi-moment method [41], and the CRP reconstruction technique [42].
However, despite the aforementioned advantages, current spatiotemporally coupled two-step fourth-order methods still suffer from significant limitations. On the one hand, such methods are currently confined to explicit frameworks. As a result, their time step is severely constrained by the CFL stability condition, and this restriction in turn drastically reduces their computational efficiency when dealing with stiff problems. On the another hand, the stiffness of a problem is often predominantly manifested in its source terms. For instance, in problems involving multi-scale physics like combustion or chemically reacting flows, the source terms can introduce significant stiffness. The key advantage of the GRP solver over the conventional Riemann solver lies in its ability to provide the temporal derivatives of physical quantities, and this derivative information inherently encodes the effects of the source terms from the governing equations. However, a critical paradox emerges: this same temporal derivative data is utilized in two distinct ways: it is essential for constructing the time-interval fluxes and serves as the cornerstone for building the two-stage fourth-order explicit scheme. The former (time-interval fluxes) can successfully incorporate information from stiff source terms, whereas the latter (the explicit scheme itself) is fundamentally unsuitable for solving stiff problems due to its explicit nature. This inherent incompatibility points to a significant limitation in the current methodological framework. Inaddition, the incorporation of temporal derivatives may allows the proposed method to achieve not only higher computational accuracy but also a more compact formulation compared to existing implicit schemes. Based on the above considerations, we conclude that there is an urgent need to develop two-stage fourth-order implicit methods to address multi-scale problems such as those involving strong stiffness.
As the foundational component of this research on spatiotemporally coupled two-stage fourth-order implicit schemes, hereafter referred to as Implicit TSFO, we develop an implicit framework that attains fourth-order accuracy in two stages through the incorporation of temporal derivatives of physical quantities. Section 2 presents the detailed formulation of the scheme via rigorous mathematical derivation. This is followed in Section 3 by a thorough analysis of the stability conditions of the implicit method. Section 4 outlines the implementation procedure based on Newton iteration. In Section 5, the efficacy of the proposed scheme is demonstrated through several canonical numerical examples, supplemented by comparative studies with classical fourth-order implicit and explicit Runge-Kutta methods. Finally, Section 6 concludes the paper and suggests potential directions for future research.
2 The Construction of the Implicit TSFO Scheme
Consider the time-dependent equations
$\frac{\partial\bf u}{\partial t}=\mathcal{L}(\bf u),$ (2.1)
where we assume that the operator $\mathcal{L}$ is sufficiently smooth, meaning all required derivatives exist. We write (2.1) as
${\bf u}^{n+\theta}={\bf u}^{n}+\int_{t_{n}}^{t_{n}+\theta\Delta t}\mathcal{L}[{\bf u}(t)]dt,$ (2.2)
where ${\bf u}^{n+\theta}={\bf u}(t_{n}+\theta\Delta t)$, $\theta\in[0,1]$, denote $\mathcal{Q}(t)=\mathcal{L}[{\bf u}(t)]$, then
$\mathcal{Q}^{\prime}(t)=\mathcal{L}_{\bf u}\mathcal{L},\quad\mathcal{Q}^{\prime\prime}(t)=\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L},\quad\mathcal{Q}^{\prime\prime\prime}(t)=\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L},$ (2.3)
and
$\int_{t_{n}}^{t_{n}+\theta\Delta t}\mathcal{L}[{\bf u}(t)]dt$ (2.4)
$=\int_{t_{n}}^{t_{n}+\theta\Delta t}\mathcal{Q}(t)dt$
$=\int_{t_{n}}^{t_{n}+\theta\Delta t}\left\{\mathcal{Q}(t_{n})+\frac{(t-t_{n})}{1!}\mathcal{Q}^{\prime}(t_{n})+\frac{(t-t_{n})^{2}}{2!}\mathcal{Q}^{\prime\prime}(t_{n})+\frac{(t-t_{n})^{3}}{3!}\mathcal{Q}^{\prime\prime\prime}(t_{n})+\mathcal{O}[(t-t_{n})^{4}]\right\}dt$
$=\frac{\theta\Delta t}{1!}\mathcal{Q}(t_{n})+\frac{(\theta\Delta t)^{2}}{2!}\mathcal{Q}^{\prime}(t_{n})+\frac{(\theta\Delta t)^{3}}{3!}\mathcal{Q}^{\prime\prime}(t_{n})+\frac{(\theta\Delta t)^{4}}{4!}\mathcal{Q}^{\prime\prime\prime}(t_{n})+\mathcal{O}[(\theta\Delta t)^{5}].$
Especially for $\theta=\frac{1}{2}$ and 1, we have
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\frac{\Delta t}{2}\mathcal{L}({\bf u}^{n})+\frac{(\Delta t)^{2}}{8}[\mathcal{L}_{\bf u}\mathcal{L}]({\bf u}^{n})+\frac{(\Delta t)^{3}}{48}[\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}]({\bf u}^{n})$ (2.5)
$\quad\quad\quad\quad\quad\;\;+\frac{(\Delta t)^{4}}{384}[\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L}]({\bf u}^{n})+\mathcal{O}(\Delta t^{5}),$
and
${\bf u}^{n+1}={\bf u}^{n}+\Delta t\mathcal{L}({\bf u}^{n})+\frac{(\Delta t)^{2}}{2}[\mathcal{L}_{\bf u}\mathcal{L}]({\bf u}^{n})+\frac{(\Delta t)^{3}}{6}[\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}]({\bf u}^{n})$ (2.6)
$\quad\quad\quad\quad\;\;\;+\frac{(\Delta t)^{4}}{24}[\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L}]({\bf u}^{n})+\mathcal{O}(\Delta t^{5}).$
We construct a two-stage fourth-order implicit scheme by utilizing the values of physical quantities along with their temporal derivatives. The proposed scheme is formulated as follows:
Stage 1. We construct the following implicit scheme by utilizing the physical quantities at the time level $t_{n}$ and $t_{n+\frac{1}{2}}$ to obtain the solution at the intermediate time level ${\bf u}^{n+\frac{1}{2}}$,
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\Delta t\left[A_{1}\mathcal{L}({\bf u}^{n})+A_{2}\mathcal{L}({\bf u}^{n+\frac{1}{2}})\right]+\Delta t^{2}\left[B_{1}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})+B_{2}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})\right],$ (2.7)
where the coefficients $A_{1}$, $A_{2}$, $B_{1}$, $B_{2}$ are to be determined to achieve fourth-order accuracy in Stage 1.
Stage 2. We construct the following implicit scheme by utilizing the physical quantities at the time level $t_{n}$, $t_{n+\frac{1}{2}}$ and $t_{n+1}$ to obtain the solution at the next time level ${\bf u}^{n+1}$,
${\bf u}^{n+1}={\bf u}^{n}$ $+\Delta t\left[A_{3}\mathcal{L}({\bf u}^{n})+A_{4}\mathcal{L}({\bf u}^{n+\frac{1}{2}})+A_{5}\mathcal{L}({\bf u}^{n+1})\right]$ (2.8)
$+\Delta t^{2}\left[B_{3}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})+B_{4}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})+B_{5}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+1})\right],$
where the coefficients $A_{3}$, $A_{4}$, $A_{5}$, $B_{3}$, $B_{4}$, $B_{5}$ are to be determined to achieve fourth-order accuracy in $\bf Stage2$.
By treating ${\bf u}^{n}$, ${\bf u}^{n+\frac{1}{2}}$, and ${\bf u}^{n+1}$ as the exact solution and using Taylor series expansions, we obtain
${\bf u}^{n+\theta}={\bf u}^{n}+\frac{\theta\Delta t}{1!}\frac{\partial{\bf u}}{\partial t}{\Big\lvert}_{t_{n}}+\frac{(\theta\Delta t)^{2}}{2!}\frac{\partial^{2}{\bf u}}{\partial t^{2}}{\Big\lvert}_{t_{n}}+\frac{(\theta\Delta t)^{3}}{3!}\frac{\partial^{3}{\bf u}}{\partial t^{3}}{\Big\lvert}_{t_{n}}+\mathcal{O}[(\theta\Delta t)^{4}],$ (2.9)
where
$\frac{\partial\bf u}{\partial t}=\mathcal{L},\quad\frac{\partial^{2}\bf u}{\partial t^{2}}=\mathcal{L}_{\bf u}\mathcal{L},\quad\frac{\partial^{3}\bf u}{\partial t^{3}}=\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}.$ (2.10)
Especially for $\theta=\frac{1}{2}$ and 1, we have
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\frac{\Delta t}{2}\mathcal{L}({\bf u}^{n})+\frac{(\Delta t)^{2}}{8}[\mathcal{L}_{u}\mathcal{L}]({\bf u}^{n})+\frac{(\Delta t)^{3}}{48}[\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}]({\bf u}^{n})+\mathcal{O}[(\Delta t)^{4}],$ (2.11)
and
${\bf u}^{n+1}={\bf u}^{n}+\Delta t\mathcal{L}({\bf u}^{n})+\frac{(\Delta t)^{2}}{2}[\mathcal{L}_{u}\mathcal{L}]({\bf u}^{n})\\ +\frac{(\Delta t)^{3}}{6}[\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}]({\bf u}^{n})+\mathcal{O}[(\Delta t)^{4}].$ (2.12)
Taylor-expand the operators $\mathcal{L}$ with respect to $\bf u$, yielding
$\mathcal{L}({\bf u}^{n+\theta})$ $=\mathcal{L}[{\bf u}^{n}+({\bf u}^{n+\theta}-{\bf u}^{n})]$ (2.13)
$=\mathcal{L}({\bf u}^{n})+\frac{{\bf u}^{n+\theta}-{\bf u}^{n}}{1!}\mathcal{L}_{\bf u}({\bf u}^{n})+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{2}}{2!}\mathcal{L}_{\bf uu}({\bf u}^{n})$
$\quad\quad\quad\quad+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{3}}{3!}\mathcal{L}_{\bf uuu}({\bf u}^{n})+\mathcal{O}[({\bf u}^{n+\theta}-{\bf u}^{n})^{4}]$
$=\mathcal{L}({\bf u}^{n})+({\bf u}^{n+\theta}-{\bf u}^{n})\mathcal{L}_{\bf u}({\bf u}^{n})+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{2}}{2}\mathcal{L}_{\bf uu}({\bf u}^{n})$
$\quad\quad\quad\quad+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{3}}{6}\mathcal{L}_{\bf uuu}({\bf u}^{n})+\mathcal{O}[({\bf u}^{n+\theta}-{\bf u}^{n})^{4}].$
Denote $\mathcal{G}=\frac{\partial}{\partial t}\mathcal{L}$, then
$\mathcal{G}=\mathcal{L}_{\bf u}\mathcal{L},\quad\mathcal{G}_{\bf u}=\mathcal{L}_{\bf uu}\mathcal{L}+\mathcal{L}_{\bf u}^{2},\quad\mathcal{G}_{\bf uu}=\mathcal{L}_{\bf uuu}\mathcal{L}+3\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}.$ (2.14)
Taylor-expand the operators $\mathcal{G}$ with respect to $\bf u$, yielding
$\mathcal{G}({\bf u}^{n+\theta})$ $=\mathcal{G}[{\bf u}^{n}+({\bf u}^{n+\theta}-{\bf u}^{n})]$ (2.15)
$=\mathcal{G}({\bf u}^{n})+\frac{{\bf u}^{n+\theta}-{\bf u}^{n}}{1!}\mathcal{G}_{\bf u}({\bf u}^{n})+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{2}}{2!}\mathcal{G}_{\bf uu}({\bf u}^{n})+\mathcal{O}[({\bf u}^{n+\theta}-{\bf u}^{n})^{3}]$
$=[\mathcal{L}_{\bf u}\mathcal{L}]({\bf u}^{n})+({\bf u}^{n+\theta}-{\bf u}^{n})[\mathcal{L}_{\bf uu}\mathcal{L}+\mathcal{L}_{\bf u}^{2}]({\bf u}^{n})$
$\quad\quad\quad\quad\quad\quad+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{2}}{2}[\mathcal{L}_{\bf uuu}\mathcal{L}+3\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}]+\mathcal{O}[({\bf u}^{n+\theta}-{\bf u}^{n})^{3}].$
Substituting (2.11), (2.13), (2.15) into (2.7), we obtain
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\Delta t\left(A_{1}+A_{2}\right)\mathcal{L}({\bf u}^{n})$ (2.16)
$\quad\quad\quad\quad\quad\;\;+(\Delta t)^{2}\left(\frac{1}{2}A_{2}+B_{1}+B_{2}\right)[\mathcal{L}_{\bf u}\mathcal{L}]({\bf u}^{n})$
$\quad\quad\quad\quad\quad\;\;+(\Delta t)^{3}\left(\frac{1}{8}A_{2}+\frac{1}{2}B_{2}\right)[\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}]({\bf u}^{n})$
$\quad\quad\quad\quad\quad\;\;+(\Delta t)^{4}\left(\frac{1}{48}A_{2}+\frac{1}{8}B_{2}\right)[\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L}]({\bf u}^{n})$
$\quad\quad\quad\quad\quad\;\;+\mathcal{O}[(\Delta t)^{5}].$
Compare (2.5) with (2.16), we obtain that the coefficients must satisfy the following equation relationship,
$\begin{cases}A_{1}+A_{2}=\frac{1}{2},\\ \frac{1}{2}A_{2}+B_{1}+B_{2}=\frac{1}{8},\\ \frac{1}{8}A_{2}+\frac{1}{2}B_{2}=\frac{1}{48},\\ \frac{1}{48}A_{2}+\frac{1}{8}B_{2}=\frac{1}{384},\end{cases}$ (2.17)
Solve the above system of equations, we obtain the unique solution
$A_{1}=\frac{1}{4},\quad A_{2}=\frac{1}{4},\quad B_{1}=\frac{1}{48},\quad B_{2}=-\frac{1}{48}.$ (2.18)
Substituting (2.11), (2.12), (2.13) and (2.15) into (2.8), we obtain
${\bf u}^{n+1}={\bf u}^{n}$ $+\Delta t\left(A_{3}+A_{4}+A_{5}\right)\mathcal{L}({\bf u}^{n})$ (2.19)
$+\Delta t^{2}\left(\frac{1}{2}A_{4}+A_{5}+B_{3}+B_{4}+B_{5}\right)[\mathcal{L}_{\bf u}\mathcal{L}]({\bf u}^{n})$
$+\Delta t^{3}\left(\frac{1}{8}A_{4}+\frac{1}{2}A_{5}+\frac{1}{2}B_{4}+B_{5}\right)[\mathcal{L}_{\bf u}^{2}\mathcal{L}+\mathcal{L}_{\bf uu}\mathcal{L}^{2}]({\bf u}^{n})$
$+\Delta t^{4}\left(\frac{1}{48}A_{4}+\frac{1}{6}A_{5}+\frac{1}{8}B_{4}+\frac{1}{2}B_{5}\right)[\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L}]({\bf u}^{n})$
$+\mathcal{O}(\Delta t^{5})$
The comparison of (2.6) and (2.19), gives
$\begin{cases}A_{3}+A_{4}+A_{5}=1,\\ \frac{1}{2}A_{4}+A_{5}+B_{3}+B_{4}+B_{5}=\frac{1}{2},\\ \frac{1}{8}A_{4}+\frac{1}{2}A_{5}+\frac{1}{2}B_{4}+B_{5}=\frac{1}{6},\\ \frac{1}{48}A_{4}+\frac{1}{6}A_{5}+\frac{1}{8}B_{4}+\frac{1}{2}B_{5}=\frac{1}{24},\end{cases}$ (2.20)
Solve the above system of equations, we obtain
$\begin{cases}A_{3}=\frac{1}{6}+4C+\frac{1}{2}D,\\ A_{4}=\frac{2}{3}-8C+2D,\\ A_{5}=\frac{1}{6}+4C-\frac{5}{2}D,\\ B_{3}=C,\\ B_{4}=D,\\ B_{5}=\frac{1}{2}D-C.\end{cases}$ (2.21)
where $C$, $D$ are real parameters to be determined. Substituting (2.18) and (2.21) into (2.7) and (2.8) respectively, we obtain the specific form of the desired two-stage fourth-order implicit (e.g., Implicit TSFO) scheme as follows:
Stage 1. Iteratively solve the following implicit scheme to obtain the solution at the intermediate time level ${\bf u}^{n+\frac{1}{2}}$,
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\frac{\Delta t}{4}\left[\mathcal{L}({\bf u}^{n})+\mathcal{L}({\bf u}^{n+\frac{1}{2}})\right]-\frac{\Delta t^{2}}{48}\left[\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})\right].$ (2.22)
Stage 2. Iteratively solve the following implicit scheme to obtain the solution at the next time level ${\bf u}^{n}$,
${\bf u}^{n+1}={\bf u}^{n}$ $+\Delta t\left[\left(\frac{1}{6}+4C+\frac{1}{2}D\right)\mathcal{L}({\bf u}^{n})+\left(\frac{2}{3}-8C+2D\right)\mathcal{L}({\bf u}^{n+\frac{1}{2}})+\left(\frac{1}{6}+4C-\frac{5}{2}D\right)\mathcal{L}({\bf u}^{n+1})\right]$ (2.23)
$+\Delta t^{2}\left[C\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})+D\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})+\left(\frac{1}{2}D-C\right)\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+1})\right].$
3 The Stability Analysis of the Implicit TSFO Scheme
To examine the stability of the implicit scheme (2.22)-(2.23), we consider the following model equation
$\mathcal{L}(\bf u)=\lambda\bf u,$ (3.1)
where $\lambda$ is the eigenvalue of the Jacobian matrix of the system (2.1), and ${\rm Re}(\lambda)\leq 0$. Combining (3.1) and (2.1), we obtain
$\frac{\partial}{\partial t}\mathcal{L}({\bf u})=\lambda^{2}{\bf u}.$ (3.2)
Subtitute (3.1)-(3.2) into the implicit scheme (2.22)-(2.23), we obtain
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\frac{z}{4}\left({\bf u}^{n}+{\bf u}^{n+\frac{1}{2}}\right)-\frac{z^{2}}{48}\left({\bf u}^{n+\frac{1}{2}}-{\bf u}^{n}\right),$ (3.3)
${\bf u}^{n+1}={\bf u}^{n}$ $+z\left[\left(\frac{1}{6}+4C+\frac{1}{2}D\right){\bf u}^{n}+\left(\frac{2}{3}-8C+2D\right){\bf u}^{n+\frac{1}{2}}+\left(\frac{1}{6}+4C-\frac{5}{2}D\right){\bf u}^{n+1}\right]$ (3.4)
$+z^{2}\left[C{\bf u}^{n}+D{\bf u}^{n+\frac{1}{2}}+\left(\frac{1}{2}D-C\right){\bf u}^{n+1}\right],$
where $z=\lambda\Delta t$. From (3.3), we obtain
${\bf u}^{n+\frac{1}{2}}=R(z){\bf u}^{n},$ (3.5)
where
$R(z)=\frac{1+\frac{z}{4}+\frac{z^{2}}{48}}{1-\frac{z}{4}+\frac{z^{2}}{48}}=1+\frac{24}{z}+\mathcal{O}\left(\frac{1}{z^{2}}\right).$ (3.6)
Substitute (3.5) into (3.4), we obtain
${\bf u}^{n+1}=G(z){\bf u}^{n},$ (3.7)
where
$G(z)$ $=\frac{1+z\left[\left(\frac{1}{6}+4C+\frac{1}{2}D\right)+\left(\frac{2}{3}-8C+2D\right)R(z)\right]+z^{2}\left[C+DR(z)\right]}{1-\left(\frac{1}{6}+4C-\frac{5}{2}D\right)z-\left(\frac{1}{2}D-C\right)z^{2}}$ (3.8)
$\sim\frac{C+D}{C-\frac{1}{2}D},\quad z\to\infty.$
Therefore, we have the following crucial conclusion.
Lemma 3.1.
The amplification factor $G$ satisfying
$G(\infty)=0,$ (3.9)
if and only if
$C+D=0,$ (3.10)
where $C$ and $D$ are the undetermined parameters in the Implicit TSFO scheme (2.22)-(2.23).
Under the condition (3.10), the amplification factor (3.8) reduces to
$G(z)=\frac{1+z\left[\frac{1}{6}+\frac{7}{2}C+\left(\frac{2}{3}-10C\right)R(z)\right]+Cz^{2}\left[1-R(z)\right]}{1-\left(\frac{1}{6}+\frac{13}{2}C\right)z+\frac{3}{2}Cz^{2}}.$ (3.11)
Our objective in the following is to determine the range of $C$ for which the Implicit TSFO scheme (2.22)-(2.23) is A-stable. Achieving this requires not only need condition (3.9) above but also need the following condition:
$\sup_{\text{Re}(z)\leq 0}|G(z)|\leq 1.$ (3.12)
The theoretical foundation for verifying A-stability is based on the maximum modulus principle: If $G(z)$ is analytic in the closed left half-plane $\text{Re}(z)\leq 0$, then the maximum of $\lvert G(z)\lvert$ occurs on the boundary, specifically on: the imaginary axis $z=iy,\;y\in\mathbb{R}$ and at infinity $z\to\infty$. Therefore, to achieve the A-stability of the Implicit TSFO scheme (2.22)-(2.23), we need to examine the following three conditions:
Condition 1: $G(z)$ is analytic in the closed left half-plane, i.e.,
$G\in\mathcal{H}(\mathbb{C}_{-}),$ (3.13)
where $\mathbb{C}_{-}=\{z\in\mathbb{C}:\mathrm{Re}(z)\leq 0\}$.
Condition 2: The $L_{\infty}$-norm of $G$ on the imaginary axis is bounded by 1, i.e.,
$\|G(i\cdot)\|_{L_{\infty}(\mathbb{R})}\leq 1,$ (3.14)
or equivalently,
$\sup_{y\in\mathbb{R}}|G(iy)|\leq 1.$ (3.15)
Condition 3: $G$ is bounded at infinity with magnitude not exceeding 1, i.e.,
$|G(\infty)|\leq 1.$ (3.16)
More precisely, the limit
$\lim_{|z|\to\infty,\,\mathrm{Re}(z)\leq 0}G(z)$ (3.17)
exists and its modulus is bounded by 1.
Since Condition 3 (3.16)-(3.17) is already satisfied by Lemma 3.1, we focus on analyzing Conditions 1 and Condition 2 to determine the range of $C$.
Examination of Condition 1, i.e., (3.13).
The first requirement for A-stability is that the amplification factor $G(z)$ must be analytic in the closed left half-plane $\mathbb{C}_{-}$. This is equivalent to requiring that all poles of $G(z)$ lie strictly in the right half-plane. The denominator of $G(z)$ is given by:
$D(z)=\left[1-\left(\frac{1}{6}+\frac{13}{2}C\right)z+\frac{3}{2}Cz^{2}\right]\left(1-\frac{z}{4}+\frac{z^{2}}{48}\right).$ (3.18)
The poles are the roots of $D(z)=0$. The equation
$1-\frac{z}{4}+\frac{z^{2}}{48}=0,$ (3.19)
has two complex roots
$z=6+2i\sqrt{3},\quad z=6-2i\sqrt{3},$ (3.20)
both located in the right half-plane. The roots of the equation
$1-\left(\frac{1}{6}+\frac{13}{2}C\right)z+\frac{3}{2}Cz^{2}=0$ (3.21)
are given by:
$z=\begin{cases}\frac{\frac{1}{3}+13C\pm\sqrt{\left(\frac{1}{3}+13C\right)^{2}-24C}}{6C},&C\neq 0,\\ 6,&C=0.\end{cases}$ (3.22)
So that the real part of $z$ is
$\text{Re}(z)=\begin{cases}\frac{\frac{1}{3}+13C}{6C},&C\neq 0,\\ 6,&C=0.\end{cases}$ (3.23)
Therefore, we have the following conclusion.
Lemma 3.2.
If and only if
$C\geq 0\quad\text{or}\quad C<-\frac{1}{39},$ (3.24)
we have
${\textnormal{Re}}(z)>0,$ (3.25)
i.e., all roots of the equation (3.21) lie in the right half-plane, in another words, the amplification factor $G(z)$ is analytic in $\textnormal{Re}(z)\leq 0$.
Examination of Condition 2, i.e., (3.14)-(3.15).
Substitute $z=iy$ into $R(z)$, we obtain
$R(iy)=\frac{1+i\frac{y}{4}-\frac{y^{2}}{48}}{1-i\frac{y}{4}-\frac{y^{2}}{48}}.$ (3.26)
Let
$N_{R}=1+i\frac{y}{4}-\frac{y^{2}}{48},\quad D_{R}=1-i\frac{y}{4}-\frac{y^{2}}{48}.$ (3.27)
Since $D_{R}=\overline{N_{R}}$ (complex conjugate), we have:
$\lvert R(iy)\lvert=\Big\lvert\frac{N_{R}}{D_{R}}\Big\lvert=\Big\lvert\frac{N_{R}}{\overline{N_{R}}}\Big\lvert=1.$ (3.28)
Thus
$R(iy)=e^{i\psi(y)},$ (3.29)
with phase
$\psi(y)=2\text{arg}\left(1+i\frac{y}{4}-\frac{y^{2}}{48}\right)=2\text{arctan}\left(\frac{y/4}{1-y^{2}/48}\right).$ (3.30)
Substitute $z=iy$ and (3.29) into $G(z)$, we obtain
$G(iy)=\frac{1+iy[\frac{1}{6}+\frac{7}{2}C+(\frac{2}{3}-10C)e^{i\psi}]-Cy^{2}(1-e^{i\psi})}{1-(\frac{1}{6}+\frac{13}{2}C)iy-\frac{3}{2}Cy^{2}}.$ (3.31)
Analytically proving $\lvert G(iy)\lvert\leq 1$ for all $y\in\mathbb{R}$ is complex. Instead, we employ a systematic numerical scanning algorithm to determine the A-stability conditions of the scheme. The parameter space $C\in[0.00,0.10]$ is discretized into 5,000 sampling points, while the frequency domain along the imaginary axis $y\in[10^{-8},10^{4}]$ is densely sampled using 25,000 logarithmically spaced points. For each parameter-frequency pair $(C,y)$, the corresponding amplification factor $\lvert G(iy)\lvert$ is computed. The algorithm identifies the maximum amplification factor across all frequencies and verifies its compliance with the stability criterion $\lvert G(iy)\lvert\leq 1$. The corresponding numerical results are shown in the figure 3.1. Based on the above systematic numerical scanning analysis, Lemma 3.1. and Lemma 3.2., we draw the following conclusion regarding the A-stability of the scheme.
Conclusion 3.3.
The Implicit TSFO scheme, given by equations (2.22)-(2.23), achieves A-stability under the condition
D=C,D=-C, (3.32)
with the parameter $C$ confined to the interval
$C\in[0.018824,0.045589],$ (3.33)
where the optimal parameter
$C=0.018924$ (3.34)
provides superior numerical damping characteristics.
Figure 3.1: A-stability: Valid C Range Determination
By applying the Conclusion 3.3. to the Implicit TSFO scheme (2.22)-(2.23), we derive the final form of this scheme, which is A-stable.
Stage 1. Iteratively solve the following implicit scheme to obtain the solution at the intermediate time level ${\bf u}^{n+\frac{1}{2}}$,
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\frac{\Delta t}{4}\left[\mathcal{L}({\bf u}^{n})+\mathcal{L}({\bf u}^{n+\frac{1}{2}})\right]-\frac{\Delta t^{2}}{48}\left[\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})\right].$ (3.35)
Stage 2. Iteratively solve the following implicit scheme to obtain the solution at the next time level ${\bf u}^{n}$,
${\bf u}^{n+1}={\bf u}^{n}$ $+\Delta t\left[\left(\frac{1}{6}+\frac{7}{2}C\right)\mathcal{L}({\bf u}^{n})+\left(\frac{2}{3}-10C\right)\mathcal{L}({\bf u}^{n+\frac{1}{2}})+\left(\frac{1}{6}+\frac{13}{2}C\right)\mathcal{L}({\bf u}^{n+1})\right]$ (3.36)
$+C\Delta t^{2}\left[\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})-\frac{3}{2}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+1})\right],$
where $C\in[0.018824,0.045589]$, with $C=0.018924$ being the optimal value.
4 The Implementation of the Implicit TSFO Scheme
To achieve accelerated convergence, the implicit scheme can be realized through Newton’s iterative method, where the initial guess for each stage uses the result of the two-stage fourth-order explicit scheme [34].
Stage 1. The solution at the intermediate time level ${\bf u}^{n+\frac{1}{2}}$ is computed iteratively as follows:
  1. 1.
    Initialization:
    $({\bf u}^{n+\frac{1}{2}})^{(0)}={\bf u}^{n}+\frac{\Delta t}{2}\mathcal{L}({\bf u}^{n})+\frac{\Delta t^{2}}{8}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})$ (4.1)
  2. 2.
    Iteration: For $k=0,1,2,\dots$, repeat the following steps until convergence:
    1. (a)
      Solve the linear system for $\Delta{\bf u}$:
      $J_{1}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]\Delta{\bf u}=-F_{1}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]$ (4.2)
    2. (b)
      Update the solution:
      $({\bf u}^{n+\frac{1}{2}})^{(k+1)}=({\bf u}^{n+\frac{1}{2}})^{(k)}+\Delta{\bf u}$ (4.3)
  3. 3.
    Convergence check: The iteration continues while
    $\left\lVert({\bf u}^{n+\frac{1}{2}})^{(k+1)}-({\bf u}^{n+\frac{1}{2}})^{(k)}\right\rVert\geq\mathrm{TOL}$ (4.4)
  4. 4.
    Final assignment: Once converged, set
    ${\bf u}^{n+\frac{1}{2}}=({\bf u}^{n+\frac{1}{2}})^{(k+1)}$ (4.5)
where the Jacobian $J_{1}$ and residual $F_{1}$ are defined as:
$J_{1}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]$ $=\frac{\Delta t}{4}\mathcal{L}_{\bf u}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]-\frac{\Delta t^{2}}{48}\left(\frac{\partial}{\partial t}\mathcal{L}\right)_{\bf u}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]-I,$ (4.6)
$F_{1}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]$ $={\bf u}^{n}+\frac{\Delta t}{4}\left\{\mathcal{L}{({\bf u}^{n})}+\mathcal{L}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]\right\}$
$\quad-{\frac{\Delta t^{2}}{48}}\left\{\frac{\partial}{\partial t}\mathcal{L}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})\right\}-({\bf u}^{n+\frac{1}{2}})^{(k)}.$
Stage 2. The solution at the next time level ${\bf u}^{n+1}$ is computed iteratively as follows:
  1. 1.
    Initialization:
    $({\bf u}^{n+1})^{(0)}={\bf u}^{n}+\Delta t\mathcal{L}({\bf u}^{n})+\frac{\Delta t^{2}}{6}\left[\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})+2\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})\right]$ (4.7)
  2. 2.
    Iteration: For $k=0,1,2,\dots$, repeat the following steps until convergence:
    1. (a)
      Solve the linear system for $\Delta{\bf u}$:
      $J_{2}\left[({\bf u}^{n+1})^{(k)}\right]\Delta{\bf u}=-F_{2}\left[({\bf u}^{n+1})^{(k)}\right]$ (4.8)
    2. (b)
      Update the solution:
      $({\bf u}^{n+1})^{(k+1)}=({\bf u}^{n+1})^{(k)}+\Delta{\bf u}$ (4.9)
  3. 3.
    Convergence check: The iteration continues while
    $\left\lVert\lvert({\bf u}^{n+1})^{(k+1)}-({\bf u}^{n+1})^{(k)}\right\rVert\geq\mathrm{TOL}$ (4.10)
  4. 4.
    Final assignment: Once converged, set
    ${\bf u}^{n+1}=({\bf u}^{n+1})^{(k+1)}$ (4.11)
where the Jacobian $J_{2}$ and residual $F_{2}$ are defined as:
$J_{2}\left[({\bf u}^{n+1})^{(k)}\right]=\Delta t\left(\frac{1}{6}+\frac{13}{2}C\right){\mathcal{L}}_{\bf u}\left[({\bf u}^{n+1})^{(k)}\right]-\frac{3}{2}C\Delta t^{2}\left(\frac{\partial}{\partial t}\mathcal{L}\right)_{\bf u}\left[({\bf u}^{n+1})^{(k)}\right]-I,$ (4.12)
$F_{2}\left[({\bf u}^{n+1})^{(k)}\right]$ (4.13)
$={\bf u}^{n}+\Delta t\left\{\left(\frac{1}{6}+\frac{7}{2}C\right)\mathcal{L}({\bf u}^{n})+\left(\frac{2}{3}-10C\right)\mathcal{L}({\bf u}^{n+\frac{1}{2}})+\left(\frac{1}{6}+\frac{13}{2}C\right)\mathcal{L}\left[({\bf u}^{n+1})^{(k)}\right]\right\}$
$\quad\quad\;\;\;+C\Delta t^{2}\left\{\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})-\frac{3}{2}\frac{\partial}{\partial t}\mathcal{L}\left[({\bf u}^{n+1})^{(k)}\right]\right\}-({\bf u}^{n+1})^{(k)}.$
5 Numerical tests
This section will validate the convergence order, stability, and competitiveness of the proposed two-stage fourth-order implicit (e.g., Implicit TSFO) scheme (3.35)-(3.36) by applying it to several classic stiff problems, with its fourth-order accuracy being verified through the extrapolation algorithm below,
$\text{Order}=\frac{\text{ln}[\text{Error}(\Delta t_{1})/\text{Error}(\Delta t_{2})]}{\text{ln}(\Delta t_{1}/\Delta t_{2})},$ (5.1)
especially, when $\Delta t_{1}=\Delta t$, $\Delta t_{2}=\Delta t/2$, we have
$\text{Order}_{L^{2}}(\Delta t)=\frac{\text{ln}[\text{Error}_{L^{2}}(\Delta t)/\text{Error}_{L^{2}}(\Delta t/2)]}{\text{ln}2},$ (5.2)
$\text{Order}_{L^{\infty}}(\Delta t)=\frac{\text{ln}[\text{Error}_{L^{\infty}}(\Delta t)/\text{Error}_{L^{\infty}}(\Delta t/2)]}{\text{ln}2},$ (5.3)
where
$\text{Error}_{L^{2}}(\Delta t)=\lVert{\bf u}_{n}(\Delta t)-{\bf u}_{e}\lVert_{L^{2}},\quad\text{Error}_{L^{2}}(\Delta t)=\lVert{\bf u}_{n}(\Delta t)-{\bf u}_{c}\lVert_{L^{2}},$ (5.4)
$\text{Error}_{L^{\infty}}(\Delta t)=\lVert{\bf u}_{n}(\Delta t)-{\bf u}_{e}\lVert_{L^{\infty}},\quad\text{Error}_{L^{\infty}}(\Delta t)=\lVert{\bf u}_{n}(\Delta t)-{\bf u}_{c}\lVert_{L^{\infty}},$ (5.5)
${\bf u}_{n}(\Delta t)$ is the numerical solution, ${\bf u}_{e}$ is the exact solution, and ${\bf u}_{c}$ is the reference solution generated by the classical fourth-order explicit Runge-Kutta method with a highly refined time step of $1.0\times 10^{-6}$. Additionally, a direct comparison is conducted between the proposed implicit TSFO scheme and the classical fourth-order implicit Runge-Kutta (e.g., Implicit RK4) method, focusing on stable time step size and convergence error, to verify the superiority of the former.
5.1 Linear System with Widely Separated Eigenvalues
This simple linear system,
$\begin{cases}\frac{du_{1}}{dt}=-1000u_{1}+1,\\ \frac{du_{2}}{dt}=-u_{2}+1,\end{cases}$ (5.6)
provides the most intuitive illustration of stiffness, with initial conditions: $u_{1}(0)=0,u_{2}(0)=0$ and time interval $t\in[0,10]$. The exact solution of this problem is
$\begin{cases}u_{1}(t)=0.001(1-e^{-1000t}),\\ u_{2}(t)=1-e^{-t}.\end{cases}$ (5.7)
The Jacobian matrix has eigenvalues at 1000-1000 and 1-1, representing two processes that decay at rates differing by a factor of one thousand. The component $u_{1}$ reaches its steady state almost instantaneously (by $t\approx 0.005$), while $u_{2}$ evolves slowly towards its steady state over a much longer period (until $t\approx 10$). For an explicit method, the step size is limited by the fastest process $(u_{1})$, requiring steps on the order of $1/1000$ even during the long interval where only the slow evolution of $u_{2}$ is of interest. Implicit methods excel here because their superior stability allows them to take large steps controlled by the accuracy requirement for $u_{2}$, efficiently skipping over the fast transient that has already decayed.
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
      $\tau$=1.000000E+00 6.697969115862E-08 –– 8.869206919424E-04 ––
    $\tau/2$=5.000000E-01 4.145569039958E-09 4.0140816724 6.187833932367E-04 0.5193706265
    $\tau/4$=2.500000E-01 2.584340652531E-10 4.0037019911 1.466069642268E-04 2.0774808416
    $\tau/8$=1.250000E-01 1.614125216792E-11 4.0009718441 4.619753432439E-07 8.3099220681
  $\tau/16$=6.250000E-02 1.008647055151E-12 4.0002591596 9.623737490873E-12 ************
  $\tau/32$=3.125000E-02 6.303729887940E-14 4.0000719087 6.013786673943E-13 4.0002516928
  $\tau/64$=1.562500E-02 3.939772721240E-15 4.0000214067 3.758452794720E-14 4.0000629032
$\tau/128$=7.812500E-03 2.462345918528E-16 4.0000070497 2.349007393733E-15 4.0000157245
$\tau/256$=3.906250E-03 1.538963415636E-17 4.0000026093 1.468125620726E-16 4.0000039311
$\tau/512$=1.953125E-03 9.618514177224E-19 4.0000010755 9.175778879022E-18 4.0000009828
Table 1: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Linear Systems with Widely Separated Eigenvalues".
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
      $\tau$=1.000000E+00 6.697969115862E-08 –– 8.869204367202E-04 ––
    $\tau/2$=5.000000E-01 4.145569039958E-09 4.0140816724 6.187833919430E-04 0.5193702144
    $\tau/4$=2.500000E-01 2.584340652531E-10 4.0037019911 1.466069642060E-04 2.0774808388
    $\tau/8$=1.250000E-01 1.614125216792E-11 4.0009718441 4.619753175467E-07 8.3099221482
  $\tau/16$=6.250000E-02 1.008647055151E-12 4.0002591596 9.623737490765E-12 ************
  $\tau/32$=3.125000E-02 6.303729887940E-14 4.0000719087 6.013786673943E-13 4.0002516928
  $\tau/64$=1.562500E-02 3.939772721240E-15 4.0000214067 3.758452794720E-14 4.0000629032
$\tau/128$=7.812500E-03 2.462345918528E-16 4.0000070497 2.349007393733E-15 4.0000157245
$\tau/256$=3.906250E-03 1.538963415636E-17 4.0000026093 1.468125620726E-16 4.0000039311
$\tau/512$=1.953125E-03 9.618514177224E-19 4.0000010755 9.175778879022E-18 4.0000009828
Table 2: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Linear Systems with Widely Separated Eigenvalues".
5.2 Robertson’s Chemical Kinetics Problem
The Robertson problem is a classical stiff system that models a three-component chemical reaction describing hydrogen-oxygen-nitrogen kinetics. This benchmark problem is widely used to test numerical methods for stiff ordinary differential equations. The system represents the concentrations of different chemical species and exhibits characteristics of extreme stiffness due to the presence of both fast and slow time scales, with a stiffness ratio of approximately $10^{6}$. The system is governed by the following equations:
$\begin{cases}\frac{du_{1}}{dt}=-0.04u_{1}+10^{4}u_{2}u_{3},\\ \frac{du_{2}}{dt}=0.04u_{1}-10^{4}u_{2}u_{3}-3\times 10^{7}u_{2}^{2},\\ \frac{du_{3}}{dt}=3\times 10^{7}u_{2}^{2},\end{cases}$ (5.8)
with initial conditions $u_{1}(0)=1$, $u_{2}(0)=0$, $u_{3}(0)=0$. This problem is characterized by three distinct time scales: an ultrafast process ($\approx 10^{-8}$ s, governed by $k_{3}=3\times 10^{7}$), an intermediate process ($\approx 10^{-4}$ s, governed by $k_{2}=10^{4}$), and a slow process ($\approx 25$ s, governed by $k_{1}=0.04$). Therefore, to effectively validate the accuracy of numerical methods, a terminal time of $t_{\text{end}}=10$ s is chosen. This time point captures the complete transition from fast stiff dynamics to slow evolution, which tests the method’s capability to handle extreme stiffness transitions, while also avoiding the reduced sensitivity in error evaluation that typically occurs near equilibrium.
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
        $\tau$=1.000000E-02 9.021072350964E-10 –– 3.854913505270E-09 ––
    $\tau/2$=5.000000E-03 2.580108817658E-10 1.8058670259 7.197422436148E-10 2.4211462484
    $\tau/4$=2.500000E-03 7.867372614174E-12 5.0354061892 7.128430647411E-11 3.3358239349
    $\tau/8$=1.250000E-03 2.330574682235E-12 1.7551961699 4.221448735445E-13 7.3997024919
  $\tau/16$=6.250000E-04 4.684507049390E-14 5.6366446928 1.549576641782E-13 1.4458640822
  $\tau/32$=3.125000E-04 1.740469668267E-15 4.7503486640 1.087571543380E-14 3.8326918986
  $\tau/64$=1.562500E-04 8.972457729393E-17 4.2778296404 6.977734145232E-16 3.9622078670
$\tau/128$=7.812500E-05 5.094883117507E-18 4.1383822665 4.389746093103E-17 3.9905492308
$\tau/256$=3.906250E-05 3.031651023500E-19 4.0708734516 2.748123534058E-18 3.9976187324
$\tau/512$=1.953125E-05 1.851325455787E-20 4.0334732502 1.718639521320E-19 3.9991079765
Table 3: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Robertson’s Chemical Kinetics Problem".
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
        $\tau$=1.000000E-02 6.379113507692E-10 –– 2.725942952427E-09 ––
    $\tau/2$=5.000000E-03 1.824484536839E-10 1.8058670259 5.089547175647E-10 2.4211461618
    $\tau/4$=2.500000E-03 5.563292362660E-12 5.0354061892 5.040760743533E-11 3.3358239189
    $\tau/8$=1.250000E-03 1.648030284841E-12 1.7551961699 2.985132390872E-13 7.3997027526
  $\tau/16$=6.250000E-04 3.312577599763E-14 5.6366446928 1.095759488245E-13 1.4458637453
  $\tau/32$=3.125000E-04 1.230746538617E-15 4.7503486640 7.690596265044E-15 3.8326919039
  $\tau/64$=1.562500E-04 6.344736420581E-17 4.2778296405 4.934198254849E-16 3.9622078681
$\tau/128$=7.812500E-05 3.602768766448E-18 4.1383822669 3.104141983081E-17 3.9905492311
$\tau/256$=3.906250E-05 2.143785699906E-19 4.0708734580 1.943293632771E-18 3.9976187332
$\tau/512$=1.953125E-05 1.309136414687E-20 4.0334733543 1.215309709126E-19 3.9991079877
Table 4: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Robertson’s Chemical Kinetics Problem".
5.3 Ozone Decomposition Reaction Problem
The ozone decomposition reaction system is a typical stiff system, characterized by the simultaneous presence of extremely short-lived, rapidly changing radical intermediates (such as oxygen atoms) and slowly varying major species (such as ozone and oxygen), resulting in significant disparities in the kinetic rates of different components. This problem is described by the following eight-dimensional system:
$\begin{cases}\frac{du_{1}}{dt}=-1.71u_{1}+0.43u_{2}+8.32u_{3}+0.0007,\\ \frac{du_{2}}{dt}=1.71u_{1}-8.75u_{2},\\ \frac{du_{3}}{dt}=-10.03u_{3}+0.43u_{4}+0.035u_{5},\\ \frac{du_{4}}{dt}=8.32u_{2}+1.71u_{3}-1.12u_{4},\\ \frac{du_{5}}{dt}=-1.745u_{5}+0.43u_{6}+0.43u_{7},\\ \frac{du_{6}}{dt}=-280u_{6}u_{8}+0.69u_{4}+1.71u_{5}-0.43u_{6}+0.69u_{7},\\ \frac{du_{7}}{dt}=280u_{6}u_{8}-1.81u_{7},\\ \frac{du_{8}}{dt}=-280u_{6}u_{8}+1.81u_{7},\end{cases}$ (5.9)
with initial conditions: $u_{1}(0)=1,u_{2}(0)=0,u_{3}(0)=0,u_{4}(0)=0,u_{5}(0)=0,u_{6}(0)=0,u_{7}(0),u_{8}(0)=0.0057$. In the numerical testing of the ozone decomposition reaction, three critical time points were selected to comprehensively evaluate the performance of the numerical method across different kinetic regimes. These points at 1.0 s, 10.0 s, and 321.8122 s respectively assess the algorithm’s capability in handling stiffness during rapid transients, convergence accuracy in the fully developed system phase, and steady-state precision after long-term integration.
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
       $\tau$=5.0000000E-01 7.304648134165E-05 –– 2.103486775801E-03 ––
    $\tau/2$=2.5000000E-01 4.368777006510E-06 4.0635134059 5.402253376233E-05 5.2830776313
    $\tau/4$=1.2500000E-01 2.757893576099E-07 3.9855907777 2.740485668104E-06 4.3010578117
    $\tau/8$=6.2500000E-02 1.726194186584E-08 3.9979001126 1.695651923233E-07 4.0145196352
  $\tau/16$=3.1250000E-02 1.079170495833E-09 3.9996000520 1.059477506743E-08 4.0004151855
  $\tau/32$=1.5625000E-02 6.745110142817E-11 3.9999369994 6.621554154568E-10 4.0000392748
  $\tau/64$=7.8125000E-03 4.215681636995E-12 4.0000041759 4.138452076597E-11 4.0000067177
$\tau/128$=3.9062500E-03 2.634780405708E-13 4.0000112892 2.586529864712E-12 4.0000014966
$\tau/256$=1.9531250E-03 1.646728646603E-14 4.0000079786 1.616580758824E-13 4.0000003629
$\tau/512$=9.7656250E-04 1.029202139129E-15 4.0000045767 1.010362911214E-14 4.0000000900
Table 5: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=1.0$.
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
       $\tau$=5.0000000E-01 5.257460149977E-05 –– 1.660009216431E-03 ––
    $\tau/2$=2.5000000E-01 3.259751779961E-06 4.0115319928 2.959645560584E-05 5.8096230285
    $\tau/4$=1.2500000E-01 2.062223567606E-07 3.9824894616 1.927179143601E-06 3.9408618221
    $\tau/8$=6.2500000E-02 1.293209504014E-08 3.9951728248 1.228186682972E-07 3.9718889151
  $\tau/16$=3.1250000E-02 8.089656191792E-10 3.9987338152 7.708323927517E-09 3.9939688558
  $\tau/32$=1.5625000E-02 5.057178443494E-11 3.9996737996 4.822552904827E-10 3.9985482291
  $\tau/64$=7.8125000E-03 3.160918871717E-12 3.9999167726 3.014846794572E-11 3.9996404695
$\tau/128$=3.9062500E-03 1.975603318786E-13 3.9999788049 1.884396367645E-12 3.9999103293
$\tau/256$=1.9531250E-03 1.234756723741E-14 3.9999945675 1.177766019856E-13 3.9999775955
$\tau/512$=9.7656250E-04 7.717237103705E-16 3.9999985829 7.361066198361E-15 3.9999943997
Table 6: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=1.0$.
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
       $\tau$=5.0000000E-01 1.675891248573E-06 –– 8.943303121509E-06 ––
    $\tau/2$=2.5000000E-01 3.642515346050E-07 2.2019215776 1.390690585893E-06 2.6850063041
    $\tau/4$=1.2500000E-01 1.746476074086E-08 4.3824162667 8.930627055194E-08 3.9608961845
    $\tau/8$=6.2500000E-02 1.088293464093E-09 4.0043073341 5.748372386454E-09 3.9575360469
  $\tau/16$=3.1250000E-02 6.835681226013E-11 3.9928387090 3.619035911141E-10 3.9894761958
  $\tau/32$=1.5625000E-02 4.283237534322E-12 3.9963115285 2.265979384053E-11 3.9973987828
  $\tau/64$=7.8125000E-03 2.680473612197E-13 3.9981418489 1.416873790744E-12 3.9993515758
$\tau/128$=3.9062500E-03 1.676383727647E-14 3.9990636048 8.856455551941E-14 3.9998380121
$\tau/256$=1.9531250E-03 1.048081474569E-15 3.9995296457 5.535440071186E-15 3.9999595104
$\tau/512$=9.7656250E-04 6.551579640740E-17 3.9997642671 3.459674317438E-16 3.9999898781
Table 7: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=10.0$.
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
       $\tau$=5.0000000E-01 1.627111303401E-06 –– 5.211344689802E-06 ––
    $\tau/2$=2.5000000E-01 3.536219744673E-07 2.2020331094 1.349971096619E-06 1.9487271620
    $\tau/4$=1.2500000E-01 1.695488884906E-08 4.3824346966 8.669408192396E-08 3.9608511959
    $\tau/8$=6.2500000E-02 1.056521292583E-09 4.0043075785 5.580256084178E-09 3.9575302772
  $\tau/16$=3.1250000E-02 6.636118877321E-11 3.9928383022 3.513197338642E-10 3.9894748094
  $\tau/32$=1.5625000E-02 4.158192376043E-12 3.9963113203 2.199711418803E-11 3.9973984426
  $\tau/64$=7.8125000E-03 2.602219852764E-13 3.9981417443 1.375437773761E-12 3.9993514911
$\tau/128$=3.9062500E-03 1.627443432415E-14 3.9990635523 8.597451492040E-14 3.9998379910
$\tau/256$=1.9531250E-03 1.017483834694E-15 3.9995296193 5.373558010232E-15 3.9999595051
$\tau/512$=9.7656250E-04 6.360313199880E-17 3.9997642539 3.358497322562E-16 3.9999898767
Table 8: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=10.0$.
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
       $\tau$=5.0048554E-01 5.844310485027E-07 –– 3.774174108398E-08 ––
    $\tau/2$=2.5004833E-01 1.231981946167E-07 2.2435357849 4.679736139609E-07 ************
    $\tau/4$=1.2502416E-01 5.899164826275E-09 4.3843265848 3.004562111915E-08 3.9612005363
    $\tau/8$=6.2512082E-02 3.675895285249E-10 4.0043431451 1.934439284412E-09 3.9571673931
  $\tau/16$=3.1253006E-02 2.308023317497E-11 3.9928065125 1.217479412149E-10 3.9893871934
  $\tau/32$=1.5625744E-02 1.445943072441E-12 3.9962951902 7.621612240772E-12 3.9973773289
  $\tau/64$=7.8126824E-03 9.047964046463E-14 3.9981337125 4.765203882270E-13 3.9993462870
$\tau/128$=3.9062938E-03 5.658390449927E-15 3.9990595278 2.978444948245E-14 3.9998366978
$\tau/256$=1.9531350E-03 3.537566339472E-16 3.9995276037 1.861535568270E-15 3.9999591827
$\tau/512$=9.7656455E-04 2.211314986480E-17 3.9997632452 1.163453836558E-16 3.9999897963
Table 9: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=321.8122$.
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
       $\tau$=5.0048554E-01 5.329967588582E-07 –– 3.531284647674E-08 ––
    $\tau/2$=2.5004833E-01 1.123553013200E-07 2.2435426293 4.267215439844E-07 ************
    $\tau/4$=1.2502416E-01 5.379973849055E-09 4.3843252272 2.739727677181E-08 3.9611905453
    $\tau/8$=6.2512082E-02 3.352376459368E-10 4.0043430876 1.763937593967E-09 3.9571610736
  $\tau/16$=3.1253006E-02 2.104892124693E-11 3.9928065125 1.110171783565E-10 3.9893857120
  $\tau/32$=1.5625744E-02 1.318684330103E-12 3.9962951920 6.949851378237E-12 3.9973769682
  $\tau/64$=7.8126824E-03 8.251644637160E-14 3.9981337135 4.345204108855E-13 3.9993461974
$\tau/128$=3.9062938E-03 5.160390443115E-15 3.9990595283 2.715928163278E-14 3.9998366754
$\tau/256$=1.9531350E-03 3.226221959847E-16 3.9995276039 1.697461925345E-15 3.9999591771
$\tau/512$=9.7656455E-04 2.016695175240E-17 3.9997632453 1.060908330213E-16 3.9999897949
Table 10: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=321.8122$.
5.4 Van der Pol Oscillator Problem
The van der Pol oscillator is a fundamental model in nonlinear dynamics, recognized for its characteristic stable limit cycle, an isolated periodic orbit that attracts neighboring trajectories. Governed by the nonlinear damping term $\nu(1-u^{2})\frac{du}{dt}$, the system exhibits self-excited oscillations whose behavior is critically influenced by the parameter $\nu$. At small values ($\nu<1$), the oscillator displays quasi-harmonic motion; as $\nu$ increases into the intermediate range ($1\leq\nu\leq 10$), distinct fast and slow time scales begin to emerge. For large $\nu$ values such as $\nu>10$, the system demonstrates pronounced stiffness, evolving into a relaxation oscillator characterized by slow energy accumulation followed by rapid discharge. In this study, we examine the strongly stiff regime with $\nu=100$, where the system exhibits extreme stiffness ($\nu\gg 100$). The model is formally expressed as the second-order equation
$\frac{d^{2}u}{dt^{2}}-\nu(1-u^{2})\frac{du}{dt}+u=0,$ (5.10)
which is commonly rewritten as a first-order system for numerical treatment by introducing $u_{1}=u$ and $u_{2}=\frac{du}{dt}$, yielding
$\begin{cases}\frac{du_{1}}{dt}=u_{2},\\ \frac{du_{2}}{dt}=\nu(1-u_{1}^{2})u_{2}-u_{1},\end{cases}$ (5.11)
with typical initial conditions $u_{1}(0)=2$, $u_{2}(0)=0$ over the time interval $t\in[0,100]$.
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
        $\tau$=1.000000E-02 4.239641324385E-03 –– 3.224176703077E-03 ––
    $\tau/2$=5.000000E-03 1.333897566871E-05 8.3121505242 2.278940261915E-04 3.8224958012
    $\tau/4$=2.500000E-03 1.687551993730E-06 2.9826440252 1.574642176302E-05 3.8552671765
    $\tau/8$=1.250000E-03 1.061075442398E-07 3.9913328120 1.006550615658E-06 3.9675323991
  $\tau/16$=6.250000E-04 6.641293212617E-09 3.9979192304 6.326789505003E-08 3.9918023140
  $\tau/32$=3.125000E-04 4.152332733106E-10 3.9994702359 3.959878774008E-09 3.9979454275
  $\tau/64$=1.562500E-04 2.595462410976E-11 3.9998585548 2.475806070187E-10 3.9994860471
$\tau/128$=7.812500E-05 1.622207101622E-12 3.9999616736 1.547516403152E-11 3.9998717060
$\tau/256$=3.906250E-05 1.013864965219E-13 4.0000205949 9.672169552167E-13 3.9999713563
$\tau/512$=1.953125E-05 6.334338803941E-15 4.0005276704 6.044905950897E-14 4.0000477364
Table 11: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Van der Pol Oscillator Problem".
Method Implicit TSFO Implicit RK4
$\Delta t$
${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
        $\tau$=1.000000E-02 4.239531568484E-03 –– 3.224093052535E-03 ––
    $\tau/2$=5.000000E-03 1.333862718107E-05 8.3121508669 2.278880751083E-04 3.8224960444
    $\tau/4$=2.500000E-03 1.687507904449E-06 2.9826440262 1.574601038175E-05 3.8552671938
    $\tau/8$=1.250000E-03 1.061047720459E-07 3.9913328121 1.006524318300E-06 3.9675324003
  $\tau/16$=6.250000E-04 6.641119700385E-09 3.9979192304 6.326624209600E-08 3.9918023141
  $\tau/32$=3.125000E-04 4.152224248130E-10 3.9994702359 3.959775317138E-09 3.9979454275
  $\tau/64$=1.562500E-04 2.595394601219E-11 3.9998585548 2.475741386603E-10 3.9994860471
$\tau/128$=7.812500E-05 1.622164719398E-12 3.9999616736 1.547475972317E-11 3.9998717060
$\tau/256$=3.906250E-05 1.013838476707E-13 4.0000205949 9.671916854432E-13 3.9999713563
$\tau/512$=1.953125E-05 6.334173311281E-15 4.0005276704 6.044748020038E-14 4.0000477364
Table 12: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Van der Pol Oscillator Problem".
Numerical results confirm that the proposed two-step fourth-order implicit (Implicit TSFO) scheme effectively solves stiff problems using large time steps without compromising fourth-order accuracy. When compared to the classical implicit Runge-Kutta (Implicit RK4) method, our scheme reduces the numerical error by about an order of magnitude under identical time steps and offers superior stability that enables even larger step sizes.
6 Conclusions and future work
This paper presents a novel two-stage fourth-order implicit scheme. We rigorously establish its fourth-order accuracy through mathematical analysis and systematically examine the scheme’s stability, deriving the conditions under which A-stability can be achieved. A detailed implementation procedure based on Newton iteration is provided. The effectiveness of the proposed method is validated through numerical experiments on several classical strongly stiff benchmark cases. Comparisons with classical explicit and implicit fourth-order Runge-Kutta methods demonstrate that the proposed scheme possesses superior stability regions and smaller numerical errors.
In future work, we will combine this implicit scheme with its explicit counterpart—a spatiotemporally coupled two-stage fourth-order method—to solve fluid dynamics problems characterized by strongly stiff source terms, such as those encountered in chemical reactions. Compared to other types of implicit schemes, the proposed method incorporates temporal derivatives of physical quantities, which is expected to enhance both computational accuracy and compactness of the scheme.
7 Acknowledgments
Zhixin Huo’s research work has been supported by the Key Program of Henan Higher Education Institutions (Grant No. 26A110007), the Young Talents Fund of Henan Province (Grant No. 252300423500), the "Double First-Class" Project of the School of Geomatics, Henan Polytechnic University (Grant No. BSJJ202306), and the Doctoral Startup Foundation of Henan Polytechnic University (Grant No. B2024-60).