### v2

# An Implicit Two-Stage Fourth-Order Temporal Discretization Scheme of Lax-Wendroff Type for Hydrodynamic Problems with Stiff Source Terms Part I: Formulation, Stability Analysis, and Newton’s Iteration

Abstract
Existing two-stage fourth-order (TSFO) methods [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] are confined to explicit frameworks and can only be applied to homogeneous hydrodynamic problems describing idealized states. More importantly, the ability of these methods to achieve fourth-order time accuracy in merely two stages relies on utilizing temporal derivatives of physical quantities, which are obtained through generalized Riemann problem (GRP) solvers. While the advantage of the GRP solvers [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27] over traditional Riemann solvers [28, 29] lies precisely in their ability to naturally incorporate the effects of source terms. Specifically, by employing the Lax-Wendroff methodology, the temporal derivatives of physical quantities are converted into spatial derivatives of fluxes and source terms through the governing equations, thereby avoiding errors arising from operator splitting. Therefore, to extend the applicability of TSFO methods to hydrodynamic problems with stiff source terms that describe real physical phenomena, and to fully leverage the advantages of these GRP solvers [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], there is an urgent need to extending TSFO methods from explicit frameworks to implicit ones. The present work lays a solid theoretical foundation and provides a concrete implementation pathway for this research direction. Firstly, this paper derives, through rigorous mathematical analysis, an implicit temporal discretization scheme that achieves fourth-order accuracy in just two stages. Subsequently, sufficient conditions for A‑stability of the proposed implicit scheme are established via systematic theoretical analysis, and a corresponding Newton iteration procedure is formulated to accelerate convergence. Finally, the proposed implicit scheme is extensively validated using classical stiff benchmark problems, confirming its effectiveness and competitiveness. Numerical results demonstrate that the proposed scheme indeed achieves fourth-order time accuracy in two stages and, compared to the classical fourth‑order implicit Runge‑Kutta method, offers a larger stable time step and smaller convergence error.
Key words: two-stage fourth-order (TSFO); generalized Riemann problem (GRP); Lax-Wendroff methodology; implicit scheme; stiff source term; A-stability
1 Introduction
Hydrodynamic problems with stiff source terms establish a fundamental mathematical framework for describing complex physical phenomena in the real world. This framework distinguishes physical processes according to their characteristic time scales: the hydrodynamic equations govern slow, wave-dominated macroscopic dynamics, while stiff source terms capture fast, localized microscopic interactions. Classical examples include shock-reaction coupling in detonation waves, pollutant-photochemical reactions in atmospheric transport, shock-radiation coupling in astrophysical flows, and the co-evolution of magnetohydrodynamic waves with atomic-scale processes in fusion plasmas. By contrast, homogeneous hydrodynamic equations reduce such complex flow systems in real physical environments to purely convective systems characterized by a single time scale. In numerical implementation, the latter, which involves only one time scale, can typically be solved efficiently using explicit schemes. The former, however, contains multiple distinct fast and slow time scales within the same system. If a purely explicit scheme is applied, the time step constrained by the CFL stability condition will be limited by the smallest time scale, leading to a sharp decline in computational efficiency. Therefore, such systems often require implicit or semi-implicit schemes, in which slowly varying fluid convection is treated explicitly while rapidly varying stiff source terms are handled implicitly, enabling efficient and stable numerical simulation across multiple scales.
The Lax-Wendroff method, serving as a discrete implementation of the Cauchy-Kovalevskaya theorem, establishes and popularizes the fundamental paradigm of directly coupling temporal and spatial derivatives through the governing equations to achieve high-order discretization. It inherently supports the embedding of stiff source terms. Its core concept can be metaphorically described as predicting temporal evolution from spatial structures. Specifically, the method employs a Taylor expansion in time and, by means of the governing equations, transforms higher-order temporal derivatives into combinations of spatial derivatives that incorporate both flux terms and source terms. As a result, temporal evolution information from source terms is automatically integrated into the construction of local time-advancement operators, thereby enabling a unified description of multiscale physical processes and avoiding numerical errors introduced by operator-splitting methods that separately treat fluxes and sources. This characteristic makes the Lax-Wendroff framework particularly suitable for developing implicit high-order schemes, offering an effective mathematical framework for solving complex multiscale flow problems in real-world physics.
The philosophical principle of Lax-Wendroff, which couples space and time through the governing equations themselves, has become a cornerstone for developing advanced numerical methods. One of its key applications is the construction of high‑accuracy generalized Riemann problem (GRP) solvers. For instance, Men’shov’s modified GRP solver [14] enhances the resolution of complex wave interactions; ADER solver [15, 16, 17, 18, 19, 20, 21] enables arbitrary‑order accuracy in both space and time; and Ben-Artiz’s nonlinear GRP solver [22, 23, 24, 25, 26, 27] provides precise treatment of strong discontinuities in compressible flows. These modern descendants are capable of handling hydrodynamic systems with stiff source terms within a unified, unsplit framework, thereby avoiding operator‑splitting errors and preserving the asymptotic properties of complex multiscale physical systems. Furthermore, such GRP solvers are essential for simulating real‑world flow fields containing shocks, turbulence, material interfaces, and intricate physical phenomena, where the regularity of the flow is typically very weak. This stems from the following reasons. Classical Riemann solvers like HLL solver [28] and Roe solver [29] are limited to resolving instantaneous states to construct Cauchy-type instantaneous fluxes, often leading to accumulated numerical errors near strong discontinuities. In contrast, GRP solvers incorporate temporal derivative information to capture local evolutionary trends, enabling construction of temporal-interval fluxes. According to finite-volume principles, these temporal-interval fluxes compensate for spatial regularity deficiencies through their built-in temporal evolution, allowing accurate resolution of strong discontinuities [30, 31].
The philosophy of the Lax-Wendroff method is applicable not only to the development of high‑accuracy GRP solvers, but also to the construction of high‑order numerical schemes. The classical Lax‑Wendroff procedure can achieve arbitrarily high‑order temporal accuracy in a single step, but it requires repeated differentiation of the governing equations. In contrast, Runge‑Kutta methods [32, 33, 34, 35] achieve high‑order temporal accuracy without differentiating the governing equations, yet they involve multi‑stage time integration, which inevitably enlarges the computational stencil. To effectively balance these two strategies, Li et al. [1] proposed a two‑stage fourth‑order (TSFO) temporal discretization scheme that makes full use of the temporal derivative information of physical quantities and attains fourth‑order accuracy in only two stages. The research team later systematically elaborated the advantages of such kind of methods [3]: (1) Computational efficiency: Under comparable conditions, its computational efficiency is twice that of a four‑stage, fourth-order Runge‑Kutta method [4]; (2) Compactness: Because the number of time‑advancement stages is halved, the size of the computational stencil is correspondingly reduced; GRP-solver-based spatiotemporally coupled HWENO‑type reconstruction techniques [5, 6] further enhance the compactness of the scheme; (3) Stability: Its stability region has been proven to be larger than that of Runge‑Kutta methods [7]; (4) Compatibility: The method can be flexibly combined with various other numerical techniques, such as the GKS solver [8], the multi‑moment method [10], and the CRP reconstruction technique [11].
Nevertheless, despite these advantages, existing TSFO methods [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] remain confined to explicit frameworks. As a result, they are only applicable to homogeneous hydrodynamic systems and cannot effectively address problems involving stiff source terms. What is more important, the capability of such schemes to achieve fourth-order temporal accuracy in merely two stages originates from their design, which fully exploits temporal derivative information of the physical quantities, which are derived from Lax-Wendroff-based GRP solvers. And the key advantage of these solvers lies precisely in their ability to integrate spatial derivatives of fluxes and source terms simultaneously into the time derivatives of physical quantities through the governing equations, all within a unified framework, thereby eliminating numerical errors that typically arise from operator splitting. Consequently, there is a clear and pressing need to develop an implicit TSFO temporal discretization scheme. The development of such a scheme would not only extend the applicability of TSFO methods from homogeneous hydrodynamic problems to systems with stiff source terms, but would also fully harness the Lax-Wendroff methodology’s inherent capacity for seamless stiff-source integration.
As the theoretical foundation and innovative core of this study, the paper is structured as follows. Chapter 2 introduces a rigorously derived implicit temporal discretization scheme that achieves fourth-order accuracy in only two stages by fully utilizing the temporal derivative information of the physical quantities. Chapter 3 establishes sufficient conditions for the A-stability of the proposed scheme through systematic theoretical analysis. Chapter 4 further develops a Newton‑iteration formulation of the scheme to accelerate convergence. Chapter 5 validates the accuracy and stability of the method using a series of classical stiff benchmark problems, and confirms its effectiveness and competitiveness by comparing its performance with classical implicit and explicit Runge‑Kutta schemes. Finally, Chapter 6 summarizes the work and outlines potential directions for future research.
2 Formulation for the Implicit TSFO Scheme
Consider the time-dependent equations
$\frac{\partial\bf u}{\partial t}=\mathcal{L}(\bf u),$ (2.1)
where we assume that the operator $\mathcal{L}$ is sufficiently smooth, meaning all required derivatives exist.
In the follows, a two-stage fourth-order implicit scheme for system (2.1) is derived through rigorous mathematical analysis. (2.1) is rewritten as
${\bf u}^{n+\theta}={\bf u}^{n}+\int_{t_{n}}^{t_{n}+\theta\Delta t}\mathcal{L}[{\bf u}(t)]dt,$ (2.2)
where ${\bf u}^{n+\theta}={\bf u}(t_{n}+\theta\Delta t)$, $\theta\in[0,1]$. Denote $\mathcal{Q}(t)=\mathcal{L}[{\bf u}(t)]$, then
$\mathcal{Q}^{\prime}(t)=\mathcal{L}_{\bf u}\mathcal{L},\quad\mathcal{Q}^{\prime\prime}(t)=\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L},\quad\mathcal{Q}^{\prime\prime\prime}(t)=\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L},$ (2.3)
and
$\int_{t_{n}}^{t_{n}+\theta\Delta t}\mathcal{L}[{\bf u}(t)]dt$ (2.4)
$=\int_{t_{n}}^{t_{n}+\theta\Delta t}\mathcal{Q}(t)dt$
$=\int_{t_{n}}^{t_{n}+\theta\Delta t}\left\{\mathcal{Q}(t_{n})+\frac{(t-t_{n})}{1!}\mathcal{Q}^{\prime}(t_{n})+\frac{(t-t_{n})^{2}}{2!}\mathcal{Q}^{\prime\prime}(t_{n})+\frac{(t-t_{n})^{3}}{3!}\mathcal{Q}^{\prime\prime\prime}(t_{n})+\mathcal{O}[(t-t_{n})^{4}]\right\}dt$
$=\frac{\theta\Delta t}{1!}\mathcal{Q}(t_{n})+\frac{(\theta\Delta t)^{2}}{2!}\mathcal{Q}^{\prime}(t_{n})+\frac{(\theta\Delta t)^{3}}{3!}\mathcal{Q}^{\prime\prime}(t_{n})+\frac{(\theta\Delta t)^{4}}{4!}\mathcal{Q}^{\prime\prime\prime}(t_{n})+\mathcal{O}[(\theta\Delta t)^{5}].$
Especially for $\theta=\frac{1}{2}$ and 1, we have
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\frac{\Delta t}{2}\mathcal{L}({\bf u}^{n})+\frac{(\Delta t)^{2}}{8}[\mathcal{L}_{\bf u}\mathcal{L}]({\bf u}^{n})+\frac{(\Delta t)^{3}}{48}[\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}]({\bf u}^{n})$ (2.5)
$\quad\quad\quad\quad\quad+\frac{(\Delta t)^{4}}{384}[\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L}]({\bf u}^{n})+\mathcal{O}(\Delta t^{5}),$
and
${\bf u}^{n+1}={\bf u}^{n}+\Delta t\mathcal{L}({\bf u}^{n})+\frac{(\Delta t)^{2}}{2}[\mathcal{L}_{\bf u}\mathcal{L}]({\bf u}^{n})+\frac{(\Delta t)^{3}}{6}[\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}]({\bf u}^{n})$ (2.6)
$\quad\quad\quad\quad\;\;\;+\frac{(\Delta t)^{4}}{24}[\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L}]({\bf u}^{n})+\mathcal{O}(\Delta t^{5}).$
The proposed two-stage fourth-order implicit scheme is given as follows:
Stage 1. The solution at the intermediate time level, $\mathbf{u}^{n+\frac{1}{2}}$, is obtained by utilizing the physical quantities and their temporal derivatives at the time levels $t_{n}$ and $t_{n+\frac{1}{2}}$,
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\Delta t\left[A_{1}\mathcal{L}({\bf u}^{n})+A_{2}\mathcal{L}({\bf u}^{n+\frac{1}{2}})\right]+\Delta t^{2}\left[B_{1}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})+B_{2}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})\right],$ (2.7)
where the coefficients $A_{1}$, $A_{2}$, $B_{1}$, $B_{2}$ are to be determined to achieve fourth-order accuracy.
Stage 2. The solution at the next time level, $\mathbf{u}^{n+1}$, is obtained by utilizing the physical quantities and their temporal derivatives at the time levels $t_{n}$, $t_{n+\frac{1}{2}}$ and $t_{n+1}$,
${\bf u}^{n+1}={\bf u}^{n}$ $+\Delta t\left[A_{3}\mathcal{L}({\bf u}^{n})+A_{4}\mathcal{L}({\bf u}^{n+\frac{1}{2}})+A_{5}\mathcal{L}({\bf u}^{n+1})\right]$ (2.8)
$+\Delta t^{2}\left[B_{3}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})+B_{4}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})+B_{5}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+1})\right],$
where the coefficients $A_{3}$, $A_{4}$, $A_{5}$, $B_{3}$, $B_{4}$, $B_{5}$ are to be determined to achieve fourth-order accuracy.
By treating ${\bf u}^{n}$, ${\bf u}^{n+\frac{1}{2}}$, and ${\bf u}^{n+1}$ as the exact solution and using Taylor series expansions, we obtain
${\bf u}^{n+\theta}={\bf u}^{n}+\frac{\theta\Delta t}{1!}\frac{\partial{\bf u}}{\partial t}{\Big\lvert}_{t_{n}}+\frac{(\theta\Delta t)^{2}}{2!}\frac{\partial^{2}{\bf u}}{\partial t^{2}}{\Big\lvert}_{t_{n}}+\frac{(\theta\Delta t)^{3}}{3!}\frac{\partial^{3}{\bf u}}{\partial t^{3}}{\Big\lvert}_{t_{n}}+\mathcal{O}[(\theta\Delta t)^{4}],$ (2.9)
where
$\frac{\partial\bf u}{\partial t}=\mathcal{L},\quad\frac{\partial^{2}\bf u}{\partial t^{2}}=\mathcal{L}_{\bf u}\mathcal{L},\quad\frac{\partial^{3}\bf u}{\partial t^{3}}=\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}.$ (2.10)
Especially for $\theta=\frac{1}{2}$ and 1, we have
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\frac{\Delta t}{2}\mathcal{L}({\bf u}^{n})+\frac{(\Delta t)^{2}}{8}[\mathcal{L}_{u}\mathcal{L}]({\bf u}^{n})+\frac{(\Delta t)^{3}}{48}[\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}]({\bf u}^{n})+\mathcal{O}[(\Delta t)^{4}],$ (2.11)
and
${\bf u}^{n+1}={\bf u}^{n}+\Delta t\mathcal{L}({\bf u}^{n})+\frac{(\Delta t)^{2}}{2}[\mathcal{L}_{u}\mathcal{L}]({\bf u}^{n})\\ +\frac{(\Delta t)^{3}}{6}[\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}]({\bf u}^{n})+\mathcal{O}[(\Delta t)^{4}].$ (2.12)
Taylor-expand the operators $\mathcal{L}$ with respect to $\bf u$, yielding
$\mathcal{L}({\bf u}^{n+\theta})$ $=\mathcal{L}[{\bf u}^{n}+({\bf u}^{n+\theta}-{\bf u}^{n})]$ (2.13)
$=\mathcal{L}({\bf u}^{n})+\frac{{\bf u}^{n+\theta}-{\bf u}^{n}}{1!}\mathcal{L}_{\bf u}({\bf u}^{n})+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{2}}{2!}\mathcal{L}_{\bf uu}({\bf u}^{n})$
$\quad\quad\quad\quad+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{3}}{3!}\mathcal{L}_{\bf uuu}({\bf u}^{n})+\mathcal{O}[({\bf u}^{n+\theta}-{\bf u}^{n})^{4}]$
$=\mathcal{L}({\bf u}^{n})+({\bf u}^{n+\theta}-{\bf u}^{n})\mathcal{L}_{\bf u}({\bf u}^{n})+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{2}}{2}\mathcal{L}_{\bf uu}({\bf u}^{n})$
$\quad\quad\quad\quad+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{3}}{6}\mathcal{L}_{\bf uuu}({\bf u}^{n})+\mathcal{O}[({\bf u}^{n+\theta}-{\bf u}^{n})^{4}].$
Denote $\mathcal{G}=\frac{\partial}{\partial t}\mathcal{L}$, then
$\mathcal{G}=\mathcal{L}_{\bf u}\mathcal{L},\quad\mathcal{G}_{\bf u}=\mathcal{L}_{\bf uu}\mathcal{L}+\mathcal{L}_{\bf u}^{2},\quad\mathcal{G}_{\bf uu}=\mathcal{L}_{\bf uuu}\mathcal{L}+3\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}.$ (2.14)
Taylor-expand the operators $\mathcal{G}$ with respect to $\bf u$, yielding
$\mathcal{G}({\bf u}^{n+\theta})$ $=\mathcal{G}[{\bf u}^{n}+({\bf u}^{n+\theta}-{\bf u}^{n})]$ (2.15)
$=\mathcal{G}({\bf u}^{n})+\frac{{\bf u}^{n+\theta}-{\bf u}^{n}}{1!}\mathcal{G}_{\bf u}({\bf u}^{n})+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{2}}{2!}\mathcal{G}_{\bf uu}({\bf u}^{n})+\mathcal{O}[({\bf u}^{n+\theta}-{\bf u}^{n})^{3}]$
$=[\mathcal{L}_{\bf u}\mathcal{L}]({\bf u}^{n})+({\bf u}^{n+\theta}-{\bf u}^{n})[\mathcal{L}_{\bf uu}\mathcal{L}+\mathcal{L}_{\bf u}^{2}]({\bf u}^{n})$
$\quad\quad\quad\quad\quad\;\;\;+\frac{({\bf u}^{n+\theta}-{\bf u}^{n})^{2}}{2}[\mathcal{L}_{\bf uuu}\mathcal{L}+3\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}]+\mathcal{O}[({\bf u}^{n+\theta}-{\bf u}^{n})^{3}].$
Substituting (2.11), (2.13), (2.15) into (2.7), we obtain
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\Delta t\left(A_{1}+A_{2}\right)\mathcal{L}({\bf u}^{n})$ (2.16)
$\quad\quad\quad\quad\;\;\;+(\Delta t)^{2}\left(\frac{1}{2}A_{2}+B_{1}+B_{2}\right)[\mathcal{L}_{\bf u}\mathcal{L}]({\bf u}^{n})$
$\quad\quad\quad\quad\;\;\;+(\Delta t)^{3}\left(\frac{1}{8}A_{2}+\frac{1}{2}B_{2}\right)[\mathcal{L}_{\bf uu}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{2}\mathcal{L}]({\bf u}^{n})$
$\quad\quad\quad\quad\;\;\;+(\Delta t)^{4}\left(\frac{1}{48}A_{2}+\frac{1}{8}B_{2}\right)[\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L}]({\bf u}^{n})$
$\quad\quad\quad\quad\;\;\;+\mathcal{O}[(\Delta t)^{5}].$
Compare (2.5) with (2.16), we obtain that the coefficients must satisfy the following system of equations,
$\begin{cases}A_{1}+A_{2}=\frac{1}{2},\\ \frac{1}{2}A_{2}+B_{1}+B_{2}=\frac{1}{8},\\ \frac{1}{8}A_{2}+\frac{1}{2}B_{2}=\frac{1}{48},\\ \frac{1}{48}A_{2}+\frac{1}{8}B_{2}=\frac{1}{384}.\end{cases}$ (2.17)
Solve the above system of equations, we obtain the unique solution,
$A_{1}=\frac{1}{4},\quad A_{2}=\frac{1}{4},\quad B_{1}=\frac{1}{48},\quad B_{2}=-\frac{1}{48}.$ (2.18)
Substituting (2.11), (2.12), (2.13) and (2.15) into (2.8), we obtain
${\bf u}^{n+1}={\bf u}^{n}$ $+\Delta t\left(A_{3}+A_{4}+A_{5}\right)\mathcal{L}({\bf u}^{n})$ (2.19)
$+\Delta t^{2}\left(\frac{1}{2}A_{4}+A_{5}+B_{3}+B_{4}+B_{5}\right)[\mathcal{L}_{\bf u}\mathcal{L}]({\bf u}^{n})$
$+\Delta t^{3}\left(\frac{1}{8}A_{4}+\frac{1}{2}A_{5}+\frac{1}{2}B_{4}+B_{5}\right)[\mathcal{L}_{\bf u}^{2}\mathcal{L}+\mathcal{L}_{\bf uu}\mathcal{L}^{2}]({\bf u}^{n})$
$+\Delta t^{4}\left(\frac{1}{48}A_{4}+\frac{1}{6}A_{5}+\frac{1}{8}B_{4}+\frac{1}{2}B_{5}\right)[\mathcal{L}_{\bf uuu}\mathcal{L}^{3}+4\mathcal{L}_{\bf uu}\mathcal{L}_{\bf u}\mathcal{L}^{2}+\mathcal{L}_{\bf u}^{3}\mathcal{L}]({\bf u}^{n})$
$+\mathcal{O}(\Delta t^{5}).$
The comparison of (2.6) and (2.19), gives
$\begin{cases}A_{3}+A_{4}+A_{5}=1,\\ \frac{1}{2}A_{4}+A_{5}+B_{3}+B_{4}+B_{5}=\frac{1}{2},\\ \frac{1}{8}A_{4}+\frac{1}{2}A_{5}+\frac{1}{2}B_{4}+B_{5}=\frac{1}{6},\\ \frac{1}{48}A_{4}+\frac{1}{6}A_{5}+\frac{1}{8}B_{4}+\frac{1}{2}B_{5}=\frac{1}{24}.\end{cases}$ (2.20)
Solve the above system of equations, we obtain
$\begin{cases}A_{3}=\frac{1}{6}+4C+\frac{1}{2}D,\\ A_{4}=\frac{2}{3}-8C+2D,\\ A_{5}=\frac{1}{6}+4C-\frac{5}{2}D,\\ B_{3}=C,\\ B_{4}=D,\\ B_{5}=\frac{1}{2}D-C,\end{cases}$ (2.21)
where $C$, $D$ are real parameters to be determined. Substituting (2.18) and (2.21) into (2.7) and (2.8) respectively, we obtain the specific form of the desired implicit two-stage fourth-order (Implicit TSFO) temporal discretization scheme as follows:
Stage 1. Iteratively solve the following implicit scheme to obtain the solution at the intermediate time level ${\bf u}^{n+\frac{1}{2}}$,
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\frac{\Delta t}{4}\left[\mathcal{L}({\bf u}^{n})+\mathcal{L}({\bf u}^{n+\frac{1}{2}})\right]-\frac{\Delta t^{2}}{48}\left[\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})\right].$ (2.22)
Stage 2. Iteratively solve the following implicit scheme to obtain the solution at the next time level ${\bf u}^{n}$,
${\bf u}^{n+1}={\bf u}^{n}$ $+\Delta t\left[\left(\frac{1}{6}+4C+\frac{1}{2}D\right)\mathcal{L}({\bf u}^{n})+\left(\frac{2}{3}-8C+2D\right)\mathcal{L}({\bf u}^{n+\frac{1}{2}})+\left(\frac{1}{6}+4C-\frac{5}{2}D\right)\mathcal{L}({\bf u}^{n+1})\right]$ (2.23)
$+\Delta t^{2}\left[C\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})+D\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})+\left(\frac{1}{2}D-C\right)\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+1})\right],$
where $C$, $D$ are real parameters to be determined.
3 Stability Analysis for the Implicit TSFO Scheme
To examine the stability of the implicit scheme (2.22)-(2.23), we consider the following model equation
$\mathcal{L}(\bf u)=\lambda\bf u,$ (3.1)
where $\lambda$ is the eigenvalue of the Jacobian matrix of the system (2.1), and ${\rm Re}(\lambda)\leq 0$. Combining (3.1) and (2.1), we obtain
$\frac{\partial}{\partial t}\mathcal{L}({\bf u})=\lambda^{2}{\bf u}.$ (3.2)
Subtitute (3.1)-(3.2) into the implicit scheme (2.22)-(2.23), we obtain
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\frac{z}{4}\left({\bf u}^{n}+{\bf u}^{n+\frac{1}{2}}\right)-\frac{z^{2}}{48}\left({\bf u}^{n+\frac{1}{2}}-{\bf u}^{n}\right),$ (3.3)
${\bf u}^{n+1}={\bf u}^{n}$ $+z\left[\left(\frac{1}{6}+4C+\frac{1}{2}D\right){\bf u}^{n}+\left(\frac{2}{3}-8C+2D\right){\bf u}^{n+\frac{1}{2}}+\left(\frac{1}{6}+4C-\frac{5}{2}D\right){\bf u}^{n+1}\right]$ (3.4)
$+z^{2}\left[C{\bf u}^{n}+D{\bf u}^{n+\frac{1}{2}}+\left(\frac{1}{2}D-C\right){\bf u}^{n+1}\right],$
where $z=\lambda\Delta t$. From (3.3), we obtain
${\bf u}^{n+\frac{1}{2}}=R(z){\bf u}^{n},$ (3.5)
where the amplification factor
$R(z)=\frac{1+\frac{z}{4}+\frac{z^{2}}{48}}{1-\frac{z}{4}+\frac{z^{2}}{48}}=1+\frac{24}{z}+\mathcal{O}\left(\frac{1}{z^{2}}\right).$ (3.6)
Substitute (3.5) into (3.4), we obtain
${\bf u}^{n+1}=G(z){\bf u}^{n},$ (3.7)
where the amplification factor
$G(z)$ $=\frac{1+z\left[\left(\frac{1}{6}+4C+\frac{1}{2}D\right)+\left(\frac{2}{3}-8C+2D\right)R(z)\right]+z^{2}\left[C+DR(z)\right]}{1-\left(\frac{1}{6}+4C-\frac{5}{2}D\right)z-\left(\frac{1}{2}D-C\right)z^{2}}$ (3.8)
$\sim\frac{C+D}{C-\frac{1}{2}D},\quad z\to\infty.$
Therefore, we have the following crucial conclusion.
Lemma 3.1.
The amplification factor $G$ satisfying
$G(\infty)=0,$ (3.9)
if and only if
$C+D=0,$ (3.10)
where $C$ and $D$ are the undetermined parameters in the Implicit TSFO scheme (2.22)-(2.23).
Under the condition (3.10), the amplification factor (3.8) reduces to
$G(z)=\frac{1+z\left[\frac{1}{6}+\frac{7}{2}C+\left(\frac{2}{3}-10C\right)R(z)\right]+Cz^{2}\left[1-R(z)\right]}{1-\left(\frac{1}{6}+\frac{13}{2}C\right)z+\frac{3}{2}Cz^{2}}.$ (3.11)
Our objective in the following is to determine the range of $C$ for which the Implicit TSFO scheme (2.22)-(2.23) is A-stable. Achieving this requires not only need condition (3.9) but also need the following condition:
$\sup_{\text{Re}(z)\leq 0}|G(z)|\leq 1.$ (3.12)
The theoretical foundation for verifying the condition (3.12) is based on the Maximum Modulus Principle: If $G(z)$ is analytic in the closed left half-plane $\text{Re}(z)\leq 0$, then the maximum of $\lvert G(z)\lvert$ occurs on the boundary, specifically on: the imaginary axis $z=iy,\;y\in\mathbb{R}$ and at infinity $z\to\infty$. Therefore, to achieve the condition (3.12) for the Implicit TSFO scheme (2.22)-(2.23), we need to examine the following three conditions:
Condition 1: $G(z)$ is analytic in the closed left half-plane, i.e.,
$G(z)\in\mathcal{H}(\mathbb{C}_{-}),$ (3.13)
where $\mathbb{C}_{-}=\{z\in\mathbb{C}:\mathrm{Re}(z)\leq 0\}$.
Condition 2: The $L_{\infty}$-norm of $G(z)$ on the imaginary axis is bounded by 1, i.e.,
$\|G(i\cdot)\|_{L_{\infty}(\mathbb{R})}\leq 1,$ (3.14)
or equivalently,
$\sup_{y\in\mathbb{R}}|G(iy)|\leq 1.$ (3.15)
Condition 3: $G(z)$ is bounded at infinity with magnitude not exceeding 1, i.e.,
$|G(\infty)|\leq 1.$ (3.16)
More precisely, the limit
$\lim_{|z|\to\infty,\,\mathrm{Re}(z)\leq 0}G(z)$ (3.17)
exists and its modulus is bounded by 1.
Since Condition 3 (3.16)-(3.17) is already satisfied by Lemma 3.1, we focus on analyzing Conditions 1 and Condition 2 to determine the range of $C$.
Examination of Condition 1, i.e., (3.13).
Condition 1 requires the amplification factor $G(z)$ to be analytic in the closed left half-plane $\mathbb{C}_{-}$, which is equivalent to requiring that all poles of $G(z)$ lie strictly in the right half-plane. The denominator of $G(z)$ is given by:
$D(z)=\left[1-\left(\frac{1}{6}+\frac{13}{2}C\right)z+\frac{3}{2}Cz^{2}\right]\left(1-\frac{z}{4}+\frac{z^{2}}{48}\right).$ (3.18)
The poles are the roots of $D(z)=0$. The equation
$1-\frac{z}{4}+\frac{z^{2}}{48}=0,$ (3.19)
has two complex roots
$z=6+2i\sqrt{3},\quad z=6-2i\sqrt{3},$ (3.20)
both located in the right half-plane. The roots of the equation
$1-\left(\frac{1}{6}+\frac{13}{2}C\right)z+\frac{3}{2}Cz^{2}=0$ (3.21)
are given by:
$z=\begin{cases}\frac{\frac{1}{3}+13C\pm\sqrt{\left(\frac{1}{3}+13C\right)^{2}-24C}}{6C},&C\neq 0,\\ 6,&C=0.\end{cases}$ (3.22)
So that the real part of $z$ is
$\text{Re}(z)=\begin{cases}\frac{\frac{1}{3}+13C}{6C},&C\neq 0,\\ 6,&C=0.\end{cases}$ (3.23)
Therefore, we have the following conclusion.
Lemma 3.2.
If and only if
$C\geq 0\quad\text{or}\quad C<-\frac{1}{39},$ (3.24)
we have
${\textnormal{Re}}(z)>0,$ (3.25)
i.e., all roots of the equation (3.21) lie in the right half-plane, in another words, the amplification factor $G(z)$ is analytic in the closed left half-plane.
Examination of Condition 2, i.e., (3.14)-(3.15).
Substitute $z=iy$ into $R(z)$, we obtain
$R(iy)=\frac{1+i\frac{y}{4}-\frac{y^{2}}{48}}{1-i\frac{y}{4}-\frac{y^{2}}{48}}.$ (3.26)
Let
$N_{R}=1+i\frac{y}{4}-\frac{y^{2}}{48},\quad D_{R}=1-i\frac{y}{4}-\frac{y^{2}}{48}.$ (3.27)
Since $D_{R}=\overline{N_{R}}$ (complex conjugate), we have:
$\lvert R(iy)\lvert=\Big\lvert\frac{N_{R}}{D_{R}}\Big\lvert=\Big\lvert\frac{N_{R}}{\overline{N_{R}}}\Big\lvert=1.$ (3.28)
Thus
$R(iy)=e^{i\psi(y)},$ (3.29)
with phase
$\psi(y)=2\text{arg}\left(1+i\frac{y}{4}-\frac{y^{2}}{48}\right)=2\text{arctan}\left(\frac{y/4}{1-y^{2}/48}\right).$ (3.30)
Substitute $z=iy$ and (3.29) into (3.11), we obtain
$G(iy)=\frac{1+iy[\frac{1}{6}+\frac{7}{2}C+(\frac{2}{3}-10C)e^{i\psi}]-Cy^{2}(1-e^{i\psi})}{1-(\frac{1}{6}+\frac{13}{2}C)iy-\frac{3}{2}Cy^{2}}.$ (3.31)
Analytically proving $\lvert G(iy)\lvert\leq 1$ for all $y\in\mathbb{R}$ is complex. Instead, we employ a systematic numerical scanning algorithm to determine the A-stability conditions of the scheme. The parameter space $C\in[0.00,0.10]$ is discretized into 5,000 sampling points, while the frequency domain along the imaginary axis $y\in[10^{-8},10^{4}]$ is densely sampled using 25,000 logarithmically spaced points. For each parameter-frequency pair $(C,y)$, the corresponding amplification factor $\lvert G(iy)\lvert$ is computed. The algorithm identifies the maximum amplification factor across all frequencies and verifies its compliance with the stability criterion $\lvert G(iy)\lvert\leq 1$. The corresponding numerical results are shown in the figure 3.1. Based on the above systematic numerical scanning analysis, Lemma 3.1. and Lemma 3.2., we draw the following conclusion regarding the A-stability of the scheme.
Conclusion 3.3.
The Implicit TSFO scheme, given by equations (2.22)-(2.23), achieves A-stability under the condition
D=C,D=-C, (3.32)
with the parameter $C$ confined to the interval
$C\in[0.018824,0.045589],$ (3.33)
where the optimal parameter
$C=0.018924$ (3.34)
provides superior numerical damping characteristics.
Figure 3.1: A-stability: Valid C Range Determination
By applying the Conclusion 3.3. to the Implicit TSFO scheme (2.22)-(2.23), we derive the final form of this scheme, which is A-stable.
Stage 1. Iteratively solve the following implicit scheme to obtain the solution at the intermediate time level ${\bf u}^{n+\frac{1}{2}}$,
${\bf u}^{n+\frac{1}{2}}={\bf u}^{n}+\frac{\Delta t}{4}\left[\mathcal{L}({\bf u}^{n})+\mathcal{L}({\bf u}^{n+\frac{1}{2}})\right]-\frac{\Delta t^{2}}{48}\left[\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})\right].$ (3.35)
Stage 2. Iteratively solve the following implicit scheme to obtain the solution at the next time level ${\bf u}^{n}$,
${\bf u}^{n+1}={\bf u}^{n}$ $+\Delta t\left[\left(\frac{1}{6}+\frac{7}{2}C\right)\mathcal{L}({\bf u}^{n})+\left(\frac{2}{3}-10C\right)\mathcal{L}({\bf u}^{n+\frac{1}{2}})+\left(\frac{1}{6}+\frac{13}{2}C\right)\mathcal{L}({\bf u}^{n+1})\right]$ (3.36)
$+C\Delta t^{2}\left[\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})-\frac{3}{2}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+1})\right].$
The effective range for parameter $C$ is $[0.018824,0.045589]$, within which $C=0.018924$ serves as a favorable choice.
4 Newton’s Iteration for the Implicit TSFO Scheme
To achieve accelerated convergence, the implicit TSFO scheme (3.35)-(3.36) can be realized through Newton’s iterative method, where the initial guess for each stage uses the result of the two-stage fourth-order explicit scheme [1].
Stage 1. The solution at the intermediate time level ${\bf u}^{n+\frac{1}{2}}$ is computed iteratively as follows:
  1. 1.
    Initialization:
    $({\bf u}^{n+\frac{1}{2}})^{(0)}={\bf u}^{n}+\frac{\Delta t}{2}\mathcal{L}({\bf u}^{n})+\frac{\Delta t^{2}}{8}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})$ (4.1)
  2. 2.
    Iteration: For $k=0,1,2,\dots$, repeat the following steps until convergence:
    1. (a)
      Solve the linear system for $\Delta{\bf u}$:
      $J_{1}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]\Delta{\bf u}=-F_{1}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]$ (4.2)
    2. (b)
      Update the solution:
      $({\bf u}^{n+\frac{1}{2}})^{(k+1)}=({\bf u}^{n+\frac{1}{2}})^{(k)}+\Delta{\bf u}$ (4.3)
  3. 3.
    Convergence check: The iteration continues while
    $\left\lVert({\bf u}^{n+\frac{1}{2}})^{(k+1)}-({\bf u}^{n+\frac{1}{2}})^{(k)}\right\rVert\geq\mathrm{TOL}$ (4.4)
  4. 4.
    Final assignment: Once converged, set
    ${\bf u}^{n+\frac{1}{2}}=({\bf u}^{n+\frac{1}{2}})^{(k+1)}$ (4.5)
where the Jacobian $J_{1}$ and residual $F_{1}$ are defined as:
$J_{1}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]$ $=\frac{\Delta t}{4}\mathcal{L}_{\bf u}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]-\frac{\Delta t^{2}}{48}\left(\frac{\partial}{\partial t}\mathcal{L}\right)_{\bf u}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]-I,$ (4.6)
$F_{1}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]$ $={\bf u}^{n}+\frac{\Delta t}{4}\left\{\mathcal{L}{({\bf u}^{n})}+\mathcal{L}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]\right\}$
$\quad\quad\;\;\;-{\frac{\Delta t^{2}}{48}}\left\{\frac{\partial}{\partial t}\mathcal{L}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})\right\}-({\bf u}^{n+\frac{1}{2}})^{(k)}.$
Stage 2. The solution at the next time level ${\bf u}^{n+1}$ is computed iteratively as follows:
  1. 1.
    Initialization:
    $({\bf u}^{n+1})^{(0)}={\bf u}^{n}+\Delta t\mathcal{L}({\bf u}^{n})+\frac{\Delta t^{2}}{6}\left[\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})+2\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})\right]$ (4.7)
  2. 2.
    Iteration: For $k=0,1,2,\dots$, repeat the following steps until convergence:
    1. (a)
      Solve the linear system for $\Delta{\bf u}$:
      $J_{2}\left[({\bf u}^{n+1})^{(k)}\right]\Delta{\bf u}=-F_{2}\left[({\bf u}^{n+1})^{(k)}\right]$ (4.8)
    2. (b)
      Update the solution:
      $({\bf u}^{n+1})^{(k+1)}=({\bf u}^{n+1})^{(k)}+\Delta{\bf u}$ (4.9)
  3. 3.
    Convergence check: The iteration continues while
    $\left\lVert\lvert({\bf u}^{n+1})^{(k+1)}-({\bf u}^{n+1})^{(k)}\right\rVert\geq\mathrm{TOL}$ (4.10)
  4. 4.
    Final assignment: Once converged, set
    ${\bf u}^{n+1}=({\bf u}^{n+1})^{(k+1)}$ (4.11)
where the Jacobian $J_{2}$ and residual $F_{2}$ are defined as:
$J_{2}\left[({\bf u}^{n+1})^{(k)}\right]=\Delta t\left(\frac{1}{6}+\frac{13}{2}C\right){\mathcal{L}}_{\bf u}\left[({\bf u}^{n+1})^{(k)}\right]-\frac{3}{2}C\Delta t^{2}\left(\frac{\partial}{\partial t}\mathcal{L}\right)_{\bf u}\left[({\bf u}^{n+1})^{(k)}\right]-I,$ (4.12)
$F_{2}\left[({\bf u}^{n+1})^{(k)}\right]$ (4.13)
$={\bf u}^{n}+\Delta t\left\{\left(\frac{1}{6}+\frac{7}{2}C\right)\mathcal{L}({\bf u}^{n})+\left(\frac{2}{3}-10C\right)\mathcal{L}({\bf u}^{n+\frac{1}{2}})+\left(\frac{1}{6}+\frac{13}{2}C\right)\mathcal{L}\left[({\bf u}^{n+1})^{(k)}\right]\right\}$
$\quad\quad\;\;\;+C\Delta t^{2}\left\{\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})-\frac{3}{2}\frac{\partial}{\partial t}\mathcal{L}\left[({\bf u}^{n+1})^{(k)}\right]\right\}-({\bf u}^{n+1})^{(k)}.$
5 Numerical Tests
This section will validate the convergence order, stability, and computational efficiency of the proposed implicit two-stage fourth-order (Implicit TSFO) scheme (3.35)-(3.36) by applying it to several classic stiff problems. Its fourth-order accuracy is verified using the following extrapolation algorithm:
$\text{Order}=\frac{\ln\!\bigl[\text{Error}(\Delta t_{1})/\text{Error}(\Delta t_{2})\bigr]}{\ln\!\bigl(\Delta t_{1}/\Delta t_{2}\bigr)},$ (5.1)
in particular, when $\Delta t_{1}=\Delta t$ and $\Delta t_{2}=\Delta t/2$, we have
$\text{Order}_{L^{2}}(\Delta t)=\frac{\ln\!\bigl[\text{Error}_{L^{2}}(\Delta t)/\text{Error}_{L^{2}}(\Delta t/2)\bigr]}{\ln 2},$ (5.2)
$\text{Order}_{L^{\infty}}(\Delta t)=\frac{\ln\!\bigl[\text{Error}_{L^{\infty}}(\Delta t)/\text{Error}_{L^{\infty}}(\Delta t/2)\bigr]}{\ln 2},$
where
$\text{Error}_{L^{2}}(\Delta t)$ $=\bigl\|\mathbf{u}_{n}(\Delta t)-\mathbf{u}_{e}\bigr\|_{L^{2}},$ $\text{Error}_{L^{2}}(\Delta t)$ $=\bigl\|\mathbf{u}_{n}(\Delta t)-\mathbf{u}_{c}\bigr\|_{L^{2}},$ (5.3)
$\text{Error}_{L^{\infty}}(\Delta t)$ $=\bigl\|\mathbf{u}_{n}(\Delta t)-\mathbf{u}_{e}\bigr\|_{L^{\infty}},$ $\text{Error}_{L^{\infty}}(\Delta t)$ $=\bigl\|\mathbf{u}_{n}(\Delta t)-\mathbf{u}_{c}\bigr\|_{L^{\infty}}.$
Here, $\mathbf{u}_{n}(\Delta t)$ is the numerical solution, $\mathbf{u}_{e}$ is the exact solution, and $\mathbf{u}_{c}$ is a reference solution generated by the classical fourth-order explicit Runge–Kutta method using a highly refined time step of $1.0\times 10^{-6}$. Furthermore, to demonstrate the superiority of the proposed implicit TSFO scheme, a direct comparison is conducted with the classical fourth-order implicit Runge–Kutta method (e.g., Implicit RK4), focusing on the allowable stable time step size and the corresponding convergence error.
5.1 Linear System with Widely Separated Eigenvalues
This simple linear system,
$\begin{cases}\frac{du_{1}}{dt}=-1000u_{1}+1,\\ \frac{du_{2}}{dt}=-u_{2}+1,\end{cases}$ (5.4)
provides the most intuitive illustration of stiffness, with initial conditions: $u_{1}(0)=0,u_{2}(0)=0$ and time interval $t\in[0,10]$. The exact solution of this problem is
$\begin{cases}u_{1}(t)=0.001(1-e^{-1000t}),\\ u_{2}(t)=1-e^{-t}.\end{cases}$ (5.5)
The Jacobian matrix has eigenvalues at 1000-1000 and 1-1, representing two processes that decay at rates differing by a factor of one thousand. The component $u_{1}$ reaches its steady state almost instantaneously (by $t\approx 0.005$), while $u_{2}$ evolves slowly towards its steady state over a much longer period (until $t\approx 10$). For an explicit method, the step size is limited by the fastest process $(u_{1})$, requiring steps on the order of $1/1000$ even during the long interval where only the slow evolution of $u_{2}$ is of interest. Implicit methods excel here because their superior stability allows them to take large steps controlled by the accuracy requirement for $u_{2}$, efficiently skipping over the fast transient that has already decayed.
Method Implicit TSFO Implicit RK4
$\Delta t$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
      $\tau$=1.000000E+00 6.697969115862E-08 –– 8.869206919424E-04 ––
    $\tau/2$=5.000000E-01 4.145569039958E-09 4.0140816724 6.187833932367E-04 0.5193706265
    $\tau/4$=2.500000E-01 2.584340652531E-10 4.0037019911 1.466069642268E-04 2.0774808416
    $\tau/8$=1.250000E-01 1.614125216792E-11 4.0009718441 4.619753432439E-07 8.3099220681
  $\tau/16$=6.250000E-02 1.008647055151E-12 4.0002591596 9.623737490873E-12 ************
  $\tau/32$=3.125000E-02 6.303729887940E-14 4.0000719087 6.013786673943E-13 4.0002516928
  $\tau/64$=1.562500E-02 3.939772721240E-15 4.0000214067 3.758452794720E-14 4.0000629032
$\tau/128$=7.812500E-03 2.462345918528E-16 4.0000070497 2.349007393733E-15 4.0000157245
$\tau/256$=3.906250E-03 1.538963415636E-17 4.0000026093 1.468125620726E-16 4.0000039311
$\tau/512$=1.953125E-03 9.618514177224E-19 4.0000010755 9.175778879022E-18 4.0000009828
Table 1: A Comparison of Implicit TSFO and RK4 Methods: $L^{2}$-Norm Error and Convergence Order for Linear Systems with Widely Separated Eigenvalues
Method Implicit TSFO Implicit RK4
$\Delta t$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
      $\tau$=1.000000E+00 6.697969115862E-08 –– 8.869204367202E-04 ––
    $\tau/2$=5.000000E-01 4.145569039958E-09 4.0140816724 6.187833919430E-04 0.5193702144
    $\tau/4$=2.500000E-01 2.584340652531E-10 4.0037019911 1.466069642060E-04 2.0774808388
    $\tau/8$=1.250000E-01 1.614125216792E-11 4.0009718441 4.619753175467E-07 8.3099221482
  $\tau/16$=6.250000E-02 1.008647055151E-12 4.0002591596 9.623737490765E-12 ************
  $\tau/32$=3.125000E-02 6.303729887940E-14 4.0000719087 6.013786673943E-13 4.0002516928
  $\tau/64$=1.562500E-02 3.939772721240E-15 4.0000214067 3.758452794720E-14 4.0000629032
$\tau/128$=7.812500E-03 2.462345918528E-16 4.0000070497 2.349007393733E-15 4.0000157245
$\tau/256$=3.906250E-03 1.538963415636E-17 4.0000026093 1.468125620726E-16 4.0000039311
$\tau/512$=1.953125E-03 9.618514177224E-19 4.0000010755 9.175778879022E-18 4.0000009828
Table 2: A Comparison of Implicit TSFO and RK4 Methods: $L^{\infty}$-Norm Error and Convergence Order for Linear Systems with Widely Separated Eigenvalues
5.2 Robertson’s Chemical Kinetics Problem
The Robertson problem is a classical stiff system that models a three-component chemical reaction describing hydrogen-oxygen-nitrogen kinetics. This benchmark problem is widely used to test numerical methods for stiff ordinary differential equations. The system represents the concentrations of different chemical species and exhibits characteristics of extreme stiffness due to the presence of both fast and slow time scales, with a stiffness ratio of approximately $10^{6}$. The system is governed by the following equations:
$\begin{cases}\frac{du_{1}}{dt}=-0.04u_{1}+10^{4}u_{2}u_{3},\\ \frac{du_{2}}{dt}=0.04u_{1}-10^{4}u_{2}u_{3}-3\times 10^{7}u_{2}^{2},\\ \frac{du_{3}}{dt}=3\times 10^{7}u_{2}^{2},\end{cases}$ (5.6)
with initial conditions $u_{1}(0)=1$, $u_{2}(0)=0$, $u_{3}(0)=0$. This problem is characterized by three distinct time scales: an ultrafast process ($\approx 10^{-8}$ s, governed by $k_{3}=3\times 10^{7}$), an intermediate process ($\approx 10^{-4}$ s, governed by $k_{2}=10^{4}$), and a slow process ($\approx 25$ s, governed by $k_{1}=0.04$). Therefore, to effectively validate the accuracy of numerical methods, a terminal time of $t_{\text{end}}=10$ s is chosen. This time point captures the complete transition from fast stiff dynamics to slow evolution, which tests the method’s capability to handle extreme stiffness transitions, while also avoiding the reduced sensitivity in error evaluation that typically occurs near equilibrium.
Method Implicit TSFO Implicit RK4
$\Delta t$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
        $\tau$=1.000000E-02 9.021072350964E-10 –– 3.854913505270E-09 ––
    $\tau/2$=5.000000E-03 2.580108817658E-10 1.8058670259 7.197422436148E-10 2.4211462484
    $\tau/4$=2.500000E-03 7.867372614174E-12 5.0354061892 7.128430647411E-11 3.3358239349
    $\tau/8$=1.250000E-03 2.330574682235E-12 1.7551961699 4.221448735445E-13 7.3997024919
  $\tau/16$=6.250000E-04 4.684507049390E-14 5.6366446928 1.549576641782E-13 1.4458640822
  $\tau/32$=3.125000E-04 1.740469668267E-15 4.7503486640 1.087571543380E-14 3.8326918986
  $\tau/64$=1.562500E-04 8.972457729393E-17 4.2778296404 6.977734145232E-16 3.9622078670
$\tau/128$=7.812500E-05 5.094883117507E-18 4.1383822665 4.389746093103E-17 3.9905492308
$\tau/256$=3.906250E-05 3.031651023500E-19 4.0708734516 2.748123534058E-18 3.9976187324
$\tau/512$=1.953125E-05 1.851325455787E-20 4.0334732502 1.718639521320E-19 3.9991079765
Table 3: A Comparison of Implicit TSFO and RK4 Methods: $L^{2}$-Norm Error and Convergence Order for Robertson’s Chemical Kinetics Problem
Method Implicit TSFO Implicit RK4
$\Delta t$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
        $\tau$=1.000000E-02 6.379113507692E-10 –– 2.725942952427E-09 ––
    $\tau/2$=5.000000E-03 1.824484536839E-10 1.8058670259 5.089547175647E-10 2.4211461618
    $\tau/4$=2.500000E-03 5.563292362660E-12 5.0354061892 5.040760743533E-11 3.3358239189
    $\tau/8$=1.250000E-03 1.648030284841E-12 1.7551961699 2.985132390872E-13 7.3997027526
  $\tau/16$=6.250000E-04 3.312577599763E-14 5.6366446928 1.095759488245E-13 1.4458637453
  $\tau/32$=3.125000E-04 1.230746538617E-15 4.7503486640 7.690596265044E-15 3.8326919039
  $\tau/64$=1.562500E-04 6.344736420581E-17 4.2778296405 4.934198254849E-16 3.9622078681
$\tau/128$=7.812500E-05 3.602768766448E-18 4.1383822669 3.104141983081E-17 3.9905492311
$\tau/256$=3.906250E-05 2.143785699906E-19 4.0708734580 1.943293632771E-18 3.9976187332
$\tau/512$=1.953125E-05 1.309136414687E-20 4.0334733543 1.215309709126E-19 3.9991079877
Table 4: A Comparison of Implicit TSFO and RK4 Methods: $L^{\infty}$-Norm Error and Convergence Order for Robertson’s Chemical Kinetics Problem
5.3 Ozone Decomposition Reaction Problem
The ozone decomposition reaction system is a typical stiff system, characterized by the simultaneous presence of extremely short-lived, rapidly changing radical intermediates (such as oxygen atoms) and slowly varying major species (such as ozone and oxygen), resulting in significant disparities in the kinetic rates of different components. This problem is described by the following eight-dimensional system:
$\begin{cases}\frac{du_{1}}{dt}=-1.71u_{1}+0.43u_{2}+8.32u_{3}+0.0007,\\ \frac{du_{2}}{dt}=1.71u_{1}-8.75u_{2},\\ \frac{du_{3}}{dt}=-10.03u_{3}+0.43u_{4}+0.035u_{5},\\ \frac{du_{4}}{dt}=8.32u_{2}+1.71u_{3}-1.12u_{4},\\ \frac{du_{5}}{dt}=-1.745u_{5}+0.43u_{6}+0.43u_{7},\\ \frac{du_{6}}{dt}=-280u_{6}u_{8}+0.69u_{4}+1.71u_{5}-0.43u_{6}+0.69u_{7},\\ \frac{du_{7}}{dt}=280u_{6}u_{8}-1.81u_{7},\\ \frac{du_{8}}{dt}=-280u_{6}u_{8}+1.81u_{7},\end{cases}$ (5.7)
with initial conditions: $u_{1}(0)=1,u_{2}(0)=0,u_{3}(0)=0,u_{4}(0)=0,u_{5}(0)=0,u_{6}(0)=0,u_{7}(0),u_{8}(0)=0.0057$. In the numerical testing of the ozone decomposition reaction, three critical time points were selected to comprehensively evaluate the performance of the numerical method across different kinetic regimes. These points at 1.0 s, 10.0 s, and 321.8122 s respectively assess the algorithm’s capability in handling stiffness during rapid transients, convergence accuracy in the fully developed system phase, and steady-state precision after long-term integration.
Method Implicit TSFO Implicit RK4
$\Delta t$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
       $\tau$=5.0000000E-01 7.304648134165E-05 –– 2.103486775801E-03 ––
    $\tau/2$=2.5000000E-01 4.368777006510E-06 4.0635134059 5.402253376233E-05 5.2830776313
    $\tau/4$=1.2500000E-01 2.757893576099E-07 3.9855907777 2.740485668104E-06 4.3010578117
    $\tau/8$=6.2500000E-02 1.726194186584E-08 3.9979001126 1.695651923233E-07 4.0145196352
  $\tau/16$=3.1250000E-02 1.079170495833E-09 3.9996000520 1.059477506743E-08 4.0004151855
  $\tau/32$=1.5625000E-02 6.745110142817E-11 3.9999369994 6.621554154568E-10 4.0000392748
  $\tau/64$=7.8125000E-03 4.215681636995E-12 4.0000041759 4.138452076597E-11 4.0000067177
$\tau/128$=3.9062500E-03 2.634780405708E-13 4.0000112892 2.586529864712E-12 4.0000014966
$\tau/256$=1.9531250E-03 1.646728646603E-14 4.0000079786 1.616580758824E-13 4.0000003629
$\tau/512$=9.7656250E-04 1.029202139129E-15 4.0000045767 1.010362911214E-14 4.0000000900
Table 5: A Comparison of Implicit TSFO and RK4 Methods: $L^{2}$-Norm Error and Convergence Order for Ozone Decomposition Reaction Problem at $t_{\text{end}}=1.0$
Method Implicit TSFO Implicit RK4
$\Delta t$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
       $\tau$=5.0000000E-01 5.257460149977E-05 –– 1.660009216431E-03 ––
    $\tau/2$=2.5000000E-01 3.259751779961E-06 4.0115319928 2.959645560584E-05 5.8096230285
    $\tau/4$=1.2500000E-01 2.062223567606E-07 3.9824894616 1.927179143601E-06 3.9408618221
    $\tau/8$=6.2500000E-02 1.293209504014E-08 3.9951728248 1.228186682972E-07 3.9718889151
  $\tau/16$=3.1250000E-02 8.089656191792E-10 3.9987338152 7.708323927517E-09 3.9939688558
  $\tau/32$=1.5625000E-02 5.057178443494E-11 3.9996737996 4.822552904827E-10 3.9985482291
  $\tau/64$=7.8125000E-03 3.160918871717E-12 3.9999167726 3.014846794572E-11 3.9996404695
$\tau/128$=3.9062500E-03 1.975603318786E-13 3.9999788049 1.884396367645E-12 3.9999103293
$\tau/256$=1.9531250E-03 1.234756723741E-14 3.9999945675 1.177766019856E-13 3.9999775955
$\tau/512$=9.7656250E-04 7.717237103705E-16 3.9999985829 7.361066198361E-15 3.9999943997
Table 6: A Comparison of Implicit TSFO and RK4 Methods: $L^{\infty}$-Norm Error and Convergence Order for Ozone Decomposition Reaction Problem at $t_{\text{end}}=1.0$
Method Implicit TSFO Implicit RK4
$\Delta t$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
       $\tau$=5.0000000E-01 1.675891248573E-06 –– 8.943303121509E-06 ––
    $\tau/2$=2.5000000E-01 3.642515346050E-07 2.2019215776 1.390690585893E-06 2.6850063041
    $\tau/4$=1.2500000E-01 1.746476074086E-08 4.3824162667 8.930627055194E-08 3.9608961845
    $\tau/8$=6.2500000E-02 1.088293464093E-09 4.0043073341 5.748372386454E-09 3.9575360469
  $\tau/16$=3.1250000E-02 6.835681226013E-11 3.9928387090 3.619035911141E-10 3.9894761958
  $\tau/32$=1.5625000E-02 4.283237534322E-12 3.9963115285 2.265979384053E-11 3.9973987828
  $\tau/64$=7.8125000E-03 2.680473612197E-13 3.9981418489 1.416873790744E-12 3.9993515758
$\tau/128$=3.9062500E-03 1.676383727647E-14 3.9990636048 8.856455551941E-14 3.9998380121
$\tau/256$=1.9531250E-03 1.048081474569E-15 3.9995296457 5.535440071186E-15 3.9999595104
$\tau/512$=9.7656250E-04 6.551579640740E-17 3.9997642671 3.459674317438E-16 3.9999898781
Table 7: A Comparison of Implicit TSFO and RK4 Methods: $L^{2}$-Norm Error and Convergence Order for Ozone Decomposition Reaction Problem at $t_{\text{end}}=10.0$
Method Implicit TSFO Implicit RK4
$\Delta t$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
       $\tau$=5.0000000E-01 1.627111303401E-06 –– 5.211344689802E-06 ––
    $\tau/2$=2.5000000E-01 3.536219744673E-07 2.2020331094 1.349971096619E-06 1.9487271620
    $\tau/4$=1.2500000E-01 1.695488884906E-08 4.3824346966 8.669408192396E-08 3.9608511959
    $\tau/8$=6.2500000E-02 1.056521292583E-09 4.0043075785 5.580256084178E-09 3.9575302772
  $\tau/16$=3.1250000E-02 6.636118877321E-11 3.9928383022 3.513197338642E-10 3.9894748094
  $\tau/32$=1.5625000E-02 4.158192376043E-12 3.9963113203 2.199711418803E-11 3.9973984426
  $\tau/64$=7.8125000E-03 2.602219852764E-13 3.9981417443 1.375437773761E-12 3.9993514911
$\tau/128$=3.9062500E-03 1.627443432415E-14 3.9990635523 8.597451492040E-14 3.9998379910
$\tau/256$=1.9531250E-03 1.017483834694E-15 3.9995296193 5.373558010232E-15 3.9999595051
$\tau/512$=9.7656250E-04 6.360313199880E-17 3.9997642539 3.358497322562E-16 3.9999898767
Table 8: A Comparison of Implicit TSFO and RK4 Methods: $L^{\infty}$-Norm Error and Convergence Order for Ozone Decomposition Reaction Problem at $t_{\text{end}}=10.0$
Method Implicit TSFO Implicit RK4
$\Delta t$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
       $\tau$=5.0048554E-01 5.844310485027E-07 –– 3.774174108398E-08 ––
    $\tau/2$=2.5004833E-01 1.231981946167E-07 2.2435357849 4.679736139609E-07 ************
    $\tau/4$=1.2502416E-01 5.899164826275E-09 4.3843265848 3.004562111915E-08 3.9612005363
    $\tau/8$=6.2512082E-02 3.675895285249E-10 4.0043431451 1.934439284412E-09 3.9571673931
  $\tau/16$=3.1253006E-02 2.308023317497E-11 3.9928065125 1.217479412149E-10 3.9893871934
  $\tau/32$=1.5625744E-02 1.445943072441E-12 3.9962951902 7.621612240772E-12 3.9973773289
  $\tau/64$=7.8126824E-03 9.047964046463E-14 3.9981337125 4.765203882270E-13 3.9993462870
$\tau/128$=3.9062938E-03 5.658390449927E-15 3.9990595278 2.978444948245E-14 3.9998366978
$\tau/256$=1.9531350E-03 3.537566339472E-16 3.9995276037 1.861535568270E-15 3.9999591827
$\tau/512$=9.7656455E-04 2.211314986480E-17 3.9997632452 1.163453836558E-16 3.9999897963
Table 9: A Comparison of Implicit TSFO and RK4 Methods: $L^{2}$-Norm Error and Convergence Order for Ozone Decomposition Reaction Problem at $t_{\text{end}}=321.8122$
Method Implicit TSFO Implicit RK4
$\Delta t$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
       $\tau$=5.0048554E-01 5.329967588582E-07 –– 3.531284647674E-08 ––
    $\tau/2$=2.5004833E-01 1.123553013200E-07 2.2435426293 4.267215439844E-07 ************
    $\tau/4$=1.2502416E-01 5.379973849055E-09 4.3843252272 2.739727677181E-08 3.9611905453
    $\tau/8$=6.2512082E-02 3.352376459368E-10 4.0043430876 1.763937593967E-09 3.9571610736
  $\tau/16$=3.1253006E-02 2.104892124693E-11 3.9928065125 1.110171783565E-10 3.9893857120
  $\tau/32$=1.5625744E-02 1.318684330103E-12 3.9962951920 6.949851378237E-12 3.9973769682
  $\tau/64$=7.8126824E-03 8.251644637160E-14 3.9981337135 4.345204108855E-13 3.9993461974
$\tau/128$=3.9062938E-03 5.160390443115E-15 3.9990595283 2.715928163278E-14 3.9998366754
$\tau/256$=1.9531350E-03 3.226221959847E-16 3.9995276039 1.697461925345E-15 3.9999591771
$\tau/512$=9.7656455E-04 2.016695175240E-17 3.9997632453 1.060908330213E-16 3.9999897949
Table 10: A Comparison of Implicit TSFO and RK4 Methods: $L^{\infty}$-Norm Error and Convergence Order for Ozone Decomposition Reaction Problem at $t_{\text{end}}=321.8122$
5.4 Van der Pol Oscillator Problem
The van der Pol oscillator is a fundamental model in nonlinear dynamics, recognized for its characteristic stable limit cycle, an isolated periodic orbit that attracts neighboring trajectories. Governed by the nonlinear damping term $\nu(1-u^{2})\frac{du}{dt}$, the system exhibits self-excited oscillations whose behavior is critically influenced by the parameter $\nu$. At small values ($\nu<1$), the oscillator displays quasi-harmonic motion; as $\nu$ increases into the intermediate range ($1\leq\nu\leq 10$), distinct fast and slow time scales begin to emerge. For large $\nu$ values such as $\nu>10$, the system demonstrates pronounced stiffness, evolving into a relaxation oscillator characterized by slow energy accumulation followed by rapid discharge. In this study, we examine the strongly stiff regime with $\nu=100$, where the system exhibits extreme stiffness ($\nu\gg 100$). The model is formally expressed as the second-order equation
$\frac{d^{2}u}{dt^{2}}-\nu(1-u^{2})\frac{du}{dt}+u=0,$ (5.8)
which is commonly rewritten as a first-order system for numerical treatment by introducing $u_{1}=u$ and $u_{2}=\frac{du}{dt}$, yielding
$\begin{cases}\frac{du_{1}}{dt}=u_{2},\\ \frac{du_{2}}{dt}=\nu(1-u_{1}^{2})u_{2}-u_{1},\end{cases}$ (5.9)
with typical initial conditions $u_{1}(0)=2$, $u_{2}(0)=0$ over the time interval $t\in[0,100]$.
Method Implicit TSFO Implicit RK4
$\Delta t$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
        $\tau$=1.000000E-02 4.239641324385E-03 –– 3.224176703077E-03 ––
    $\tau/2$=5.000000E-03 1.333897566871E-05 8.3121505242 2.278940261915E-04 3.8224958012
    $\tau/4$=2.500000E-03 1.687551993730E-06 2.9826440252 1.574642176302E-05 3.8552671765
    $\tau/8$=1.250000E-03 1.061075442398E-07 3.9913328120 1.006550615658E-06 3.9675323991
  $\tau/16$=6.250000E-04 6.641293212617E-09 3.9979192304 6.326789505003E-08 3.9918023140
  $\tau/32$=3.125000E-04 4.152332733106E-10 3.9994702359 3.959878774008E-09 3.9979454275
  $\tau/64$=1.562500E-04 2.595462410976E-11 3.9998585548 2.475806070187E-10 3.9994860471
$\tau/128$=7.812500E-05 1.622207101622E-12 3.9999616736 1.547516403152E-11 3.9998717060
$\tau/256$=3.906250E-05 1.013864965219E-13 4.0000205949 9.672169552167E-13 3.9999713563
$\tau/512$=1.953125E-05 6.334338803941E-15 4.0005276704 6.044905950897E-14 4.0000477364
Table 11: A Comparison of Implicit TSFO and RK4 Methods: $L^{2}$-Norm Error and Convergence Order for Van der Pol Oscillator Problem
Method Implicit TSFO Implicit RK4
$\Delta t$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
        $\tau$=1.000000E-02 4.239531568484E-03 –– 3.224093052535E-03 ––
    $\tau/2$=5.000000E-03 1.333862718107E-05 8.3121508669 2.278880751083E-04 3.8224960444
    $\tau/4$=2.500000E-03 1.687507904449E-06 2.9826440262 1.574601038175E-05 3.8552671938
    $\tau/8$=1.250000E-03 1.061047720459E-07 3.9913328121 1.006524318300E-06 3.9675324003
  $\tau/16$=6.250000E-04 6.641119700385E-09 3.9979192304 6.326624209600E-08 3.9918023141
  $\tau/32$=3.125000E-04 4.152224248130E-10 3.9994702359 3.959775317138E-09 3.9979454275
  $\tau/64$=1.562500E-04 2.595394601219E-11 3.9998585548 2.475741386603E-10 3.9994860471
$\tau/128$=7.812500E-05 1.622164719398E-12 3.9999616736 1.547475972317E-11 3.9998717060
$\tau/256$=3.906250E-05 1.013838476707E-13 4.0000205949 9.671916854432E-13 3.9999713563
$\tau/512$=1.953125E-05 6.334173311281E-15 4.0005276704 6.044748020038E-14 4.0000477364
Table 12: A Comparison of Implicit TSFO and RK4 Methods: $L^{\infty}$-Norm Error and Convergence Order for Van der Pol Oscillator Problem
The numerical experiments confirm that the proposed implicit TSFO time discretization scheme successfully achieves fourth-order temporal accuracy in only two computational stages. The results also demonstrate the robust stability of the proposed algorithm when solving stiff problems with relatively large time steps, thereby validating its effectiveness for stiff systems. In comparison to the classical fourth-order implicit Runge-Kutta method, the present scheme produces smaller errors under identical time step sizes, typically reducing the error by an order of magnitude (reference solutions are computed using the explicit fourth-order Runge-Kutta method). Moreover, it maintains stability under even larger time steps, highlighting its superior numerical stability. The source code for all test cases presented is publicly available in the accompanying GitHub repository at: https://github.com/ZhixinHuo/Implicit-TSFO.
6 Conclusions and future work
This study fully leverages the time derivatives of physical quantities to construct a implicit TSFO temporal discretization scheme through rigorous mathematical derivation. A systematic theoretical analysis provides sufficient conditions for the scheme to achieve A-stability, and a corresponding Newton iteration process is proposed to accelerate convergence. The proposed algorithm is validated on a series of classical stiff benchmark problems, demonstrating its robustness and competitiveness. Numerical experiments show that the scheme achieves fourth-order temporal accuracy using only two stages. Compared with classical fourth-order implicit Runge-Kutta methods, it not only allows larger stable time steps but also yields smaller convergence errors.
This study establishes a solid theoretical foundation and a concrete implementation pathway for extending TSFO time discretization methods from existing explicit frameworks [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] to implicit frameworks. Its significance and necessity lie not only in extending the applicability of TSFO methods from idealized homogeneous hydrodynamic problems to realistic ones with stiff source terms, but also in unlocking the inherent potential of Lax-Wendroff-based GRP solvers to naturally handle stiff source terms within the TSFO framework. The ability of TSFO methods to achieve fourth-order temporal accuracy with only two stages stems from their utilization of the time derivatives of physical quantities, which are precisely provided by Lax-Wendroff-based GRP solvers. Compared to traditional Riemann solvers, these advanced solvers can compute temporal derivatives that simultaneously integrate the physical effects of both fluxes and source terms within a unified framework through the governing equations, thereby effectively avoiding the numerical errors introduced by operator-splitting approaches.
Future work will integrate the existing explicit TSFO time discretization method [1] with the implicit TSFO scheme proposed in this paper, and further combine it with a more compact HWENO spatial discretization method based on GRP solver [5]. This integrated framework aims to solve hydrodynamic problems involving stiff source terms, targeting the simulation of complex, multi-scale, and multi-physics phenomena in real-world systems, such as chemically reacting flows, radiation hydrodynamics, combustion dynamics, and other stiff physical systems.
7 Acknowledgments
Zhixin Huo’s research work has been supported by the Key Program of Henan Higher Education Institutions (Grant No. 26A110007), the Young Talents Fund of Henan Province (Grant No. 252300423500), the Double First-Class Project of the School of Geomatics of Henan Polytechnic University (Grant No. BSJJ202306), and the Doctoral Startup Foundation of Henan Polytechnic University (Grant No. B2024-60).